### v2

# Approximate $2$-hop neighborhoods on incremental graphs: An efficient lazy approach

Abstract.
In this work, we propose, analyze and empirically validate a lazy-update approach to maintain accurate approximations of the $2$-hop neighborhoods of dynamic graphs resulting from sequences of edge insertions.
We first show that under random input sequences, our algorithm exhibits an optimal trade-off between accuracy and insertion cost: it only performs $O(\frac{1}{\varepsilon})$ (amortized) updates per edge insertion, while the estimated size of any vertex’s $2$-hop neighborhood is at most a factor $\varepsilon$ away from its true value in most cases, regardless of the underlying graph topology and for any $\varepsilon>0$.
As a further theoretical contribution, we explore adversarial scenarios that can force our approach into a worst-case behavior at any given time $t$ of interest. We show that while worst-case input sequences do exist, a necessary condition for them to occur is that the girth of the graph released up to time $t$ be at most $4$.
Finally, we conduct extensive experiments on a collection of real, incremental social networks of different sizes, which typically have low girth. Empirical results are consistent with and typically better than our theoretical analysis anticipates. This further supports the robustness of our theoretical findings: forcing our algorithm into a worst-case behavior not only requires topologies characterized by a low girth, but also carefully crafted input sequences that are unlikely to occur in practice.
Combined with standard sketching techniques, our lazy approach proves an effective and efficient tool to support key neighborhood queries on large, incremental graphs, including neighborhood size, Jaccard similarity between neighborhoods and, in general, functions of the union and/or intersection of $2$-hop neighborhoods.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/Gnumlab/graph_ball.
1. Introduction
In this paper, we consider the task of processing a possibly large, dynamic graph $G(V,E)$, incrementally provided as a stream of edge insertions, so that at any point of the stream it is possible to efficiently evaluate different queries that involve functions of the $h$-hop neighborhoods of its vertices. For a vertex $v\in V$, its $h$-hop neighborhood is simply the set of vertices that are within $h$ hops from $v$. In the remainder, $h$-hop neighborhoods are called $h$-balls for brevity. As concrete examples of query types we consider, one might want to estimate the size of the $2$-ball at any given vertex, or the Jaccard similarity between the $2$-balls centered at any given two vertices, or other indices of a similar flavor that depend on the intersection or union between $1$-balls and/or $2$-balls, just to mention a few.
Neighborhood-based indices are common in key mining tasks, such as link prediction in social (liben2003link, ) and biological networks (wang2023assessment, ) or to describe statistical properties of large social graphs (becchetti2008efficient, ). For example, $2$-hop neighborhoods are important in social network analysis and similarity-based link prediction (zhou2021experimental, ; zareie2020similarity, ; Sim-Nodes_Survey_2024, ), while accurate approximations of $h$-balls’ sizes are used to estimate key statistical properties of (very) large social networks (boldi2011hyperanf, ; backstrom2012four, ), or as link-based features in classifiers for Web spam detection (becchetti2008link, ). Further tasks that may explicitly involve or benefit from $2$-hop neighborhood queries are accurate estimates of centrality measures that are widely adopted in social network analysis (freeman2002centrality, ; rochat2009closeness, ; bergamini2019computing, ) or mining of (usually large) bipartite networks that are frequent in certain applications (sariyuce2018peeling, ; pavlopoulos2018bipartite, ). We investigate the former example in detail in Section 6.
When the graph is static, an effective approach to this general task is to treat $h$-balls as subsets of the vertices of the graph, suitably represented using approximate summaries or sketches (agarwal2013mergeable, ). This line of attack has proved successful, for example in the efficient and scalable evaluation of important neighborhood-based queries on massive graphs that in part or mostly reside on secondary storage (feigenbaum2005graph, ; mcgregor2014graph, ; becchetti2008efficient, ; boldi2011hyperanf, ).
Nowadays, standard applications in social network analysis often entail dynamic scenarios in which input graphs evolve over time, under a sequence of edge insertions and possibly deletions (aggarwal2014evolutionary, ). In many cases, such as co-authorship, citation or user-item association networks, the corresponding dynamic graphs are inherently or mostly incremental, i.e., either edges and/or vertices are added but not deleted, or insertions account for the vast majority of updates (leskovec2007graph, ; borner2004simultaneous, ). [^1] Of course, notable examples exist in which deletions are as important, such as some affiliation networks, or networks that describe on-line messaging behavior, where edges representing message exchanges become less and less important as they age.
[…]
To address the aforementioned issues for graphs that reside in main memory, one might want to trade some degree of accuracy for the following broad goals: 1) designing algorithms with low update costs, possibly $O(1)$ amortized per edge insertion; 2) minimizing memory footprint beyond what is needed to store the graph; 3) maintaining $1$- and $2$-balls using data structures that afford efficient, real-time computation of queries as the ones mentioned earlier with minimal memory footprint.
Heavy updates are natural and well-known candidates for efficient (albeit approximate) implementation using compact, sketch-based data structures (gibbons2001estimating, ; broder2001completeness, ; broder2000identifying, ; agarwal2013mergeable, ; trevisan/646978.711822, ). However, sketches alone are of no avail in handling light updates, whose sheer potential number requires a novel approach. The literature on efficient data structures that handle insertions and often deletions over dynamic graphs is rich. However, efficient solutions to implement neighborhood-based queries on dynamic edge streams are only known for $1$-balls (BSS20, ; MROS, ; VOS, ; CGPS24, ), nor do approaches devised for other dynamic problems adapt to our setting in any obvious way, something we elaborate more upon in Section 1.2.
1.1. Our Contribution
In this paper, we propose an approach that trades some degree of accuracy for a substantial improvement in the average number of light updates. In a nutshell, upon an edge insertion, our algorithm performs the (two) corresponding heavy updates, but in general only a subset of the required light updates, according to a scheme that combines a threshold-based mechanism and a randomized, batch-update policy. Hence, for every vertex $u$, we only keep an approximation (a subset to be specific) of $u$’s $2$-ball. If $1$- and $2$-balls are represented with suitable data sketches, our approach affords constant average update cost per edge insertion. [^3] The particular sketch used depends on the neighborhood queries we want to be able to serve. When sketches are used, the cost of merging two neighborhoods corresponds to the cost of combining the corresponding sketches, which is typically a constant that depends on the desired approximation guarantees. For example, if we are interested in the Jaccard similarity between pairs of $1$- and/or $2$-balls, this cost will be proportional to the (constant) number of minhash values we use to represent each neighborhood. While the behavior and accuracy guarantees of most sketching techniques are well understood, the estimation error induced by lazy updates can be arbitrarily high in some cases. The main focus of this paper is on the latter aspect, which is absent in the static case but critical in the dynamic setting. Accordingly, we assume lossless representations of $1$-balls and approximate $2$-balls in our theoretical analyses in Sections 2, 3 and 4, while we use standard sketching techniques to represent $1$- and $2$-balls in the actual implementations of the algorithms and baselines we consider in the experimental analysis discussed in Section 5.
Almost-optimal performance on random sequences.
We prove in Section 3 that even a simplified, deterministic variant of our lazy-update Algorithm 3 achieves asymptotically optimal expected performance when the sequence of edge insertions is a random, uniform permutation over an arbitrary set of edges. In other words, our lazy approach is robust to adversarial topologies as long as the edge sequence follows a random order. Formally, we prove that, for any desired $10<\varepsilon<$ 1, our algorithm only performs $O(\frac{1}{\varepsilon})$ (amortized) updates per edge insertion, while at any time $t$ and for every vertex $v$, the estimated size of $v$’s $2$-ball is, in expectation, at most a factor $\varepsilon$ away from its true value. We further prove that this approximation result holds with a probability that exponentially increases with the true size of the $2$-ball itself (Theorem 2). Thanks to this analysis in concentration, our results can be extended to other functions of $2$-balls, including union, intersection and Jaccard similarity (see Corollary 3 for this less obvious case).
As positive as this result may sound, it begs the following questions from a careful reader: 1) Are the results above robust to adversarial sequences? 2) Is a performance analysis under random sequences representative of practical scenarios? 3) More generally, does our lazy scheme offer significant practical advantages?
Performance analysis on adversarial inputs.
While our results for random sequences are optimal regardless of the underlying graph’s topology, one might wonder about the ability of an adversary to design worst-case, adaptive sequences that force our approach to behave poorly and, in this case, whether any conditions on the graph topology are necessary for this to happen. We investigate this issues in Section 4, where we first show that it is possible to design worst-case sequences of edge insertions that force our algorithm to perform arbitrarily worse than the random setting (Theorem 1). However, as a further contribution, we also prove that worst-case input sequences exist only if the girth (diestel2024graph, ) of the final graph is at most $4$. More precisely, we show that a randomized, special case of Algorithm 3 achieves asymptotically optimal performance on a class of graphs that contains all graphs with girth at least $5$,  [^4] The class is in fact more general since it also includes graphs with a “bounded” number of cycles of length at most 4. See Definition 2, for a formal definition of this class. even when the input sequence is chosen by an adaptive adversary.
Experimental analysis.
An analysis under random permutation sequences is relatively common in the literature on dynamic edge streams and data streams (buriol2006counting, ; kapralov2014approximating, ; peng2018estimating, ; Hanauer22DynamicSurvey, ). Yet, one might reasonably wonder about its practical significance for the task considered in this paper. We address these issues in Section 5, where we conduct experiments on small, medium and large-sized, incremental graphs (whose main properties are summarized in Table 1). At least on the diverse sample of real networks we consider, experimental results on the estimation of key queries such as size and Jaccard similarity are consistent with the theoretical findings from Section 3, while results on running times highlight the significance of our approach in practice: i) light updates crucially affect computational costs in practical scenarios; ii) when combined with sketches, our approach strikes an excellent balance between loss in accuracy and computational efficiency. In particular, its accuracy is comparable or only slightly worse than the baseline that performs all necessary light updates, while it achieves speed-ups that increase with network size and can be orders of magnitude with respect to the baseline on large or very large social networks (Tables 4 and 5). We finally note that our datasets are samples of real social networks. As such, they have relatively large local and global clustering coefficients [^5] At least the undirected ones. and thus low girth. Hence, our experimental analysis further supports the robustness of our theoretical findings: forcing our algorithm(s) into a worst-case behavior not only requires topologies characterized by a low girth, but also carefully crafted input sequences that are unlikely to occur in practice.
Use cases from real networks.
While neighborhood-based queries are extremely important in social network analysis, the main focus of this paper is on crucial, new challenges that arise in dynamic settings and on possible strategies to address them. For this reason, we investigate basic neighborhood queries without specific applications in mind. As simple as it may appear however, the incremental setting we consider provides a feasible framework to tackle important mining tasks over large, dynamic networks. As an example, we consider the problem of maintaining the harmonic centrality-based ranking of the vertices of an incremental graph (rochat2009closeness, ; boldi2014axioms, ), a key mining task in social network analysis (boldi2014axioms, ; freeman2002centrality, ). Exact solutions for this problem necessarily incur high computational costs in some cases, [^6] Even in the static setting (bergamini2019computing, ). while the study of dynamic strategies with update costs of practical interest is still an open area of research (HanauerHS22, ). In Section 6, we experimentally show that our approach is a very good fit for this task, affording continuous, accurate tracking of top-$k$ ranking vertices, thus highlighting its effectiveness for a key downstream task that would otherwise be hardly feasible on large, incremental graphs.
Many real networks are inherently bipartite, describing interactions between entities of two different types. Examples include many biological networks (pavlopoulos2018bipartite, ), social networks (LATAPY200831, ), user-URL or user-item interactions (leskovec2009meme, ) to name a few. Moreover, many of them are also inherently or mostly incremental in nature. [^7] Edge/vertex deletions are rare or absent. In such networks, the prediction of future interactions is a key mining task, for which (Jaccard) similarity between $2$-hop neighborhoods can prove a valuable index for prediction purposes. For example, in author-paper networks, the set $L_{2}(v)$ of vertices at distance exactly $2$ from an author $v$ is the set of $v$’s co-authors. Hence, for any two authors $u$ and $v$ with no currently co-authored paper, the Jaccard similarity between $L_{2}(u)$ and $L_{2}(v)$ can be used to estimate the probability of $u$ and $v$ co-authoring a paper in the next future. [^8] Our approach also allows estimation of the Jaccard similarity between $L_{2}(u)$ and $L_{2}(v)$.
Final remarks.
We stress that although we present them for the undirected case for ease of presentation and for the sake of simplicity, our algorithms apply to directed graphs in general, [^9] Of course, in this case we have directed $h$-balls, i.e., sets of a vertices that can be reached in at most $h$ hops from a given vertex, or from which it is possible to reach the vertex under consideration in at most $h$ hops. while our analysis extends to the directed case with minor modifications. Moreover, our algorithms seamlessly address incremental settings in which new vertices may connect to the current graph, the latter being for example the typical scenario in co-authorship networks (borner2004simultaneous, ).
1.2. Further related work
Efficient data structures for queries that involve $h$-balls of a dynamic graph turn out to be useful in different network applications, as we mentioned earlier (becchetti2008link, ; zareie2020similarity, ; Sim-Nodes_Survey_2024, ).
As remarked in the introduction, efficient solutions for Jaccard similarity queries on $1$-balls have been proposed for different dynamic graph models: all of them share the use of suitable data sketches to manage insertion and deletion of elements from sets. In particular, (BSS20, ; MROS, ; VOS, ) proposes and compares different approaches that work in the fully-dynamic streaming model, while an efficient solution, based on a buffered version of the $k$-min-hashing scheme is proposed in (CGPS24, ). This works in the fully-dynamic streaming model and allows recovery actions when certain “worst-case” edge deletion events occur. A further algorithm is presented in (zhang2022effective, ), where bottom-$k$ sketches (cohen2007summarizing, ) are used to perform dynamic graph clustering based on Jaccard similarity among vertices’ neighborhoods. We remark that none of these previous approaches include ideas or tools that can be adapted to efficiently manage the $2$-ball update-operations we need to implement in this work.
As for other queries that might be "related" or "useful" in our setting, a considerable amount of work on data structures that support edge insertions and deletions exists for several queries, such as connectivity or reachability, (exact or approximate) distances, minimum spanning tree, (approximate) betweenness centrality, and so on. We refer the reader to (HanauerHS22, ) for a nice survey on experimental and theoretical results on the topic. To the best of our knowledge however, none of these approaches can be obviously adapted to handle the types of queries we consider in this work. For example, a natural idea would be using an incremental data structure to dynamically maintain the first $h$ levels of a BFS tree, such as (EvenS81, ; RodittyZ11, ), that achieve $O(h)$ amortized update time. However, let alone effectiveness in efficiently serving queries as the ones we consider here, the data structure uses $\Omega(n)$ space per BFS. This is prohibitive in our setting, where we would need to instantiate one such data structure for each vertex, with total space $\Omega(n^{2})$. Moreover, since in a degree-$\Delta$ graph $\Theta(\Delta)$ BFS trees can change following a single edge insertion, the corresponding amortized time per edge insertion could be as high as $\Theta(\Delta)$, which is basically the same cost of the baseline solution we discuss at the beginning of Section 2.
2. Lazy-Update Algorithms
After giving some preliminaries we will use through all this paper, in Section 2.1 we describe the lazy-update algorithmic scheme, while in Section 2.2, we provide a general bound on its amortized update cost that holds for arbitrary sequences of edge insertions.
Preliminaries and notations
The dynamic (incremental) graph model we study can be defined as a sequence ${\mathcal{G}}=\{G^{(0)}(V,E^{(0)}),\ldots,$ $G^{(t)}(V,E^{(t)}),\ldots G^{(T)}(V,E^{(T)})\}$, where: (i) the set of vertices $V=\{1,\ldots,n\}$ is fixed, (ii) $T\leq\binom{n}{2}$ is the final graph, while (iii) $E^{(t)}$ is the subset of edges at time $t$. Note that this changes in every time step $t\geq 1$, as a new edge $e^{(t)}$ is inserted, so that $E^{(t+1)}=E^{(t)}\cup\{e^{(t)}\}$. We remark that our analysis and all our results can be easily adapted to a more general model that includes any combination of the following variants: (i) growing vertex sets, (ii) multiple insertions of the same edge, and (iii) directed edges (thus yielding directed graphs). However, the corresponding adaptations of our analysis would require significantly heavier notation and some technicalities that we decided to avoid for the sake of clarity.
Our goal is to design algorithms that, at every time step $t\geq 1$, are able to efficiently compute queries over the current $2$-balls of $G^{(t)}$. As mentioned in the introduction, our focus is on queries that are typical in graph mining such as: (i) given a vertex $u$, estimate the size of $B_{2}(u)$, and (ii) given two vertices $u,v\in V$, estimate the Jaccard similarity of the corresponding $2$-balls:
$\mathbf{J}(B_{2}(u),B_{2}(v))\ =\ \frac{|B_{2}(u)\cap B_{2}(u)|}{|B_{2}(u)\cup B_{2}(u)|}\,.$
Both the theoretical and experimental analysis of our lazy-update algorithms consider the following key performance measures: the amortized update time per edge insertion and the approximation ratio of our algorithms on the quantities $|B_{2}(u)|$ and $\mathbf{J}(B_{2}(u),B_{2}(v))$, for any choice of the input vertices. Intuitively, the amortized update time is the average time it takes to process a new edge, a more formal definition is deferred to Section 2, after a detailed description of the algorithms we consider.
We next summarize notation that is used in the remainder of the paper. For a vertex $v\in V$ of a graph $G(V,E)$, we define:
$\mathcal{N}(v)$::
the set of neighborhoods of the vertex $v$.
$\deg_{v}$::
the degree of $v$. Notice that $\deg_{v}=|\mathcal{N}(v)|$;
$L_{h}(v)$::
set of vertices at distance exactly $h$ from $v$;
$B_{h}(v)$::
set of vertices at distance at most $h$ from $v$.
The reader may have noticed that, in our notation above, the term $t$ does not appear: this is due to the fact that our analysis holds at any (arbitrarily fixed) time step, which is always clear from context.
2.1. Algorithm description
Consider the addition of a new edge $(u,v)$ to $G$. Clearly, the only $2$-balls that are affected are those centered at $u$, $v$, and at every vertex $w\in\mathcal{N}(u)\cup\mathcal{N}(v)$. A baseline strategy, given as Algorithm 1 for the sake of reference, tracks changes exactly and thus updates all $2$-balls that are affected by an edge insertion.
1 Function Insert($u,v$):
2       $B_{2}(u)\leftarrow B_{2}(u)\cup B_{1}(v)$
3       foreach $w\in\mathcal{N}(u)\setminus\{v\}$ do
4             $B_{2}(w)\leftarrow B_{2}(w)\cup\{v\}$
5
6       end foreach
7      $B_{2}(v)\leftarrow B_{2}(v)\cup B_{1}(u)$
8       foreach $w\in\mathcal{N}(v)\setminus\{u\}$ do
9             $B_{2}(w)\leftarrow B_{2}(w)\cup\{u\}$
10
11       end foreach
12
13 end
Algorithm 1 Baseline algorithm.
The magnitude of the changes (and the associated computational costs) induced by Insert$(u,v)$ vary. In particular, $B_{2}(u)$ can change significantly, as all vertices in $B_{1}(v)$ will be included in $B_{2}(u)$ (we refer to this as a heavy update, see Algorithm 1 at Algorithm 1). Instead, for any vertex $w\in\mathcal{N}(u)\setminus\{v\}$, $B_{2}(w)$ will grow by at most one element, namely $v$ (this is referred to as a light update, see Algorithm 1 at Algorithm 1). Symmetrically, the same holds for $v$ and for every $w\in\mathcal{N}(v)\setminus\{u\}$ (Algorithms 1 and 1).
A key observation at this point is that, while heavy updates can be addressed using (possibly, approximate) data structures that allow efficient merging of $1$- and $2$-balls, this line of attack fails with light updates, whose cost derives from their potential number, which can be large in many real cases, as we noted in the introduction.
A first idea to reduce the average number of updates per edge insertion is to perform heavy updates immediately, instead processing light updates in batches that are performed occasionally. More precisely, when a new edge $(u,v)$ arrives, it is initially marked as a red edge. Whenever the number of red edges incident to a vertex $u$ exceeds a certain threshold, all the corresponding light updates are processed, and the state of red edges is updated to black. See Figure 1 for an example.
Figure 1. Example of insertion of a new edge $(u,v)$. The algorithm merges the $1$-ball of $v$ with the $2$-ball of $u$ (heavy update), while it does not immediately add vertex $v$ to the $2$-ball of vertex $w$ or any other of $u$’s neighbors.
The idea behind the threshold-based approach is to maintain a balance between the number of black and red edges for every vertex. While useful when edge insertions appear in a random order, this approach may fail when red edges considerably expand the original size of the $2$-ball of some vertex $u$. In order to mitigate this problem, our algorithm uses a second ingredient: upon each edge insertion $(u,v)$, the algorithm selects $k$ vertices from $\mathcal{N}(u)$ and $k$ from $\mathcal{N}(v)$ uniformly at random and performs a batch of light updates for the selected vertices, even if the threshold has not been reached yet.
These ideas are formalized in Algorithm 3. For each vertex $v$, our algorithm maintains two sets $\hat{B}_{1}(v)$ and $\hat{B}_{2}(v)$, as well as the black degree $\Delta_{v}$ and red degree $\delta_{v}$ of $v$. Our algorithm guarantees that $\hat{B}_{1}(v)$ is exactly $B_{1}(v)$, while $\hat{B}_{2}(v)$ is in general a subset of $B_{2}(v)$. The algorithm uses two global parameters, namely a threshold $\varphi\in[0,1]$, and an integer $k$. The role of the parameter $\varphi$ can be understood as follows: when $\varphi$ is set to 00, the algorithm performs all heavy and light updates for every edge insertion, ensuring that $\hat{B}_{2}(v)$ always matches $B_{2}(v)$. As $\varphi$ increases, the update function becomes lazier: light updates are not always executed, and $\hat{B}_{2}(v)$ is typically a proper subset of $B_{2}(v)$. For instance, when $\varphi=1$, light updates are performed in batches every time the degree of a vertex doubles. Parameter $k$ specifies the number of neighbors of $v$ that are randomly selected for update of their $2$-balls whenever an edge insertion involving $v$ occurs. This mechanism corresponds to Algorithms 3, 3 and 3 of Algorithm 3.
We call Lazy-Alg$(\varphi,k)$ the algorithm that runs Algorithm 2 on an initial graph $G^{(0)}$ and then processes a sequence $S$ of edge insertions by running Algorithm 3 on each edge of $S$.
Data: An undirected graph $G=(V,E)$, a threshold parameter $0\leq\varphi\leq 1$, and an integer $k\geq 0$.
1 set $\varphi$ and $k$ as global parameters
2 foreach vertex $u\in V$ do
3       $\delta_{u}\leftarrow 0$
4       $\Delta_{u}\leftarrow\deg_{u}$
5       $\hat{B}_{1}(u)\leftarrow B_{1}(u)$
6       $\hat{B}_{2}(u)\leftarrow B_{2}(u)$
7
8 end foreach
Algorithm 2 Init operation
1 Function Insert($(u,v)$):
2       for $x\in\{u,v\}$ do
3             let $y\in\{u,v\}\setminus\{x\}$
4             $\hat{B}_{1}(x)\leftarrow\hat{B}_{1}(x)\cup\{y\}$
             // heavy update
5             $\hat{B}_{2}(x)\leftarrow\hat{B}_{2}(x)\cup\hat{B}_{1}(y)$
6             $\delta_{x}\leftarrow\delta_{x}+1$
7
8            if $\delta_{x}\geq\varphi\cdot\Delta_{x}$ then
9                   $\Delta_{x}\leftarrow\Delta_{x}+\delta_{x}$
10                   $\delta_{x}\leftarrow 0$
11                   foreach $z\in\mathcal{N}(x)$ do
                         // batch of light updates
12                         $\hat{B}_{2}(z)\leftarrow\hat{B}_{2}(z)\cup\hat{B}_{1}(x)$
13
14                   end foreach
15
16             else
17                   select $k$ vertices $w_{1},\dots,w_{k}\in\mathcal{N}(x)$ u.a.r.
18                   for $i=1,\dots,k$ do
                         // batch of light updates
19                         $\hat{B}_{2}(w_{i})\leftarrow\hat{B}_{2}(w_{i})\cup\hat{B}_{1}(x)$
20
21                   end for
22
23             end if
24
25       end for
26
27 end
28
Algorithm 3 Insert
A note on neighborhood representation.
As we mentioned in the introduction, we treat $\hat{B}_{1}(v)$ and $\hat{B}_{2}(v)$ as sets of vertices in this section and in Section 4. We remark that this only serves the purpose of analyzing the error introduced by our lazy update policies: lossless representations of $1$- and $2$-balls may be unfeasible for medium or large graphs and compact data sketches are typically used to represent them in such cases. The choice of the actual sketch strongly depends on the type of query (or queries) one wants to support, such as $1$- or $2$-ball sizes (flajolet1985probabilistic, ; boldi2011hyperanf, ) or Jaccard similarity between $2$-balls (broder2000identifying, ; cohen2007summarizing, ; becchetti2008efficient, ). All sketches used for typical neighborhood queries are well-understood and come with strong accuracy guarantees. Moreover, they allow to perform the union of $1$- and $2$-balls we are interested in time proportional to the sketch size, which is independent of the sizes of the balls to merge (agarwal2013mergeable, ).
2.2. Cost analysis for arbitrary sequences
Consistently to what we remarked above, our cost analysis focuses on the number of set-union operations: This performance measure in fact dominates the computational cost of Algorithm 3. More in detail, given any sequence $S$ of edge insertions, starting from an initial graph $G^{(0)}$, we denote by $T(S)$ the overall number of union operations performed in Algorithms 3, 3, 3 and 3 of Algorithm 3 on the input sequence $S$.
We observe that a trivial upper bound to $T(S)$ is $O(\Delta|S|)$, since each insertion can cost $O(\Delta)$ union operations where $\Delta$ is the maximum degree of the current graph. However, this trivial argument turns out to be too pessimistic: in what follows, we provide a more refined analysis of the amortized cost [^10] The amortized analysis is a well-known method originally introduced in (Tarjan_amortized, ) that allows to compute tight bounds on the cost of a sequence of operations, rather than the worst-case cost of an individual operation. In more detail, we average the cost of a worst case sequence of operations to obtain a more meaningful cost per operation. per edge insertion. We say that an algorithm has amortized cost $\hat{c}$ per edge insertion if, for any sequence $S$ of edge insertions, we have $T(S)\leq\hat{c}|S|$.
Lemma 0.
Given any initial graph $G^{(0)}$ and any sequence $S$ of edge insertions, the amortized update cost of Algorithm 3 is $O(\frac{1}{\varphi}+k)$ per edge insertion.
Proof.
Let us first consider the case $k=0$, i.e., when the random selection and the consequent instructions in Algorithms 3, 3 and 3 are never performed. Our amortized analysis makes use of the accounting method (Tarjan_amortized, ). The idea is paying the cost of any batch of light updates by charging it to previous edge insertions. More precisely, we assign credits to each edge insertion that we will use to pay the cost of subsequent batches of light updates. Formally, the amortized cost of an edge insertion is defined as the actual cost of the operation, plus the credits we assign to it, minus the credits (accumulated from previous operations) we spend for it. We need to carefully define such credits in order to guarantee that the sum of the amortized costs is an upper bound to the sum of the actual costs, i.e. we always have enough credits to pay for costly batch light updates.
We proceed as follows. When we insert the edge $(u,v)$, we put $2/\varphi$ credits on $u$ and $2/\varphi$ credits on $v$. Now we bound the actual and amortized cost of each insertion.
First, consider an edge insertion $(u,v)$ that does not trigger a batch of light updates. Its actual cost is 4 union operations (those in Algorithms 3 and 3, 2 for each endpoint of $(u,v)$). Then its amortized cost is upper-bounded by $4+4/\varphi=O(1/\varphi)$. Now consider the case in which the insertion causes a batch of light updates for $u$, or $v$, or both. We show that the credits accumulated by previous insertions are sufficient to pay for its cost. To see this, consider a batch of light updates involving vertex $x\in\{u,v\}$. And let $\Delta_{x}$ and $\delta_{x}$ be the current black and red degrees of $x$ at that time (just before Algorithm 3 is evaluated). It is clear that for vertex $x$ we have accumulated $\delta_{x}\cdot 2/\varphi$ credits that now we use to pay for the cost of Algorithms 3 and 3. This cost equals to $\deg_{x}$ union operations, where $\deg_{x}$ is the current degree of vertex $x$. Since the batch of light updates has just been triggered, we have that $\delta_{x}\geq\varphi\Delta_{x}$, and hence we have at least $\delta_{x}\cdot 2/\varphi\geq\varphi\Delta_{x}\cdot 2/\varphi=2\Delta_{x}$ credits to pay for the $\deg_{x}=\Delta_{x}+\delta_{x}=\Delta_{x}+\varphi\Delta_{x}\leq 2\Delta_{x}$ union operations. This concludes the proof.
Finally, to obtain the claim when $k>0$, we notice that, in this case, every edge insertion causes $O(k)$ additional union operations. ∎
3. Random edge sequences
In this section, we analyze the accuracy of our lazy-update algorithm(s) over an arbitrary dynamic graph, whose edges are given in input as a uniformly sampled, random permutation over its edge set. Dynamic graphs resulting from random sequences of edge insertions have been an effective tool to provide theoretical insights that have often proved robust to empirical validation in various dynamic scenarios (monemizadeh2017testable, ; kapralov2014approximating, ; peng2018estimating, ; mcgregor2014graph, ; chakrabarti2008robust, ). In more detail, assume $G=(V,E)$, with $|E|=t$, is the graph observed up to some time $t$ of interest. Following (monemizadeh2017testable, ; peng2018estimating, ), we assume that the sequence of edges up to time $t$ is chosen uniformly at random from the set of all permutations over $E$ [^11] It should be noted that this includes the general case in which $t$ is any intermediate point of a longer stream that possibly extends well beyond $t$. In this case, it is well-known and easy to see that, conditioned on the subset $E$ of the edges released up to time $t$, their sequence is just a permutation of $E$. The following fact is an immediate consequence of well-known and intuitive properties of random permutations. We state it informally for the sake of completeness, avoiding any further, unnecessary notation.
Fact 1.
Consider a dynamic graph $G=(V,E)$, whose edges are observed sequentially according to a permutation over $E$ chosen uniformly at random. Then, for every $E^{\prime}\subseteq E$, the sequence in which edges in $E^{\prime}$ are observed is itself a uniformly chosen, random permutation over $E^{\prime}$.
In the remainder, for an arbitrary vertex $v$, we analyze how well the output $\hat{B}_{2}(v)$ of Algorithm 3 approximates $B_{2}(v)$ at any round $t$ in terms of its coverage:
Definition 3.0.
We say that the output $\hat{B}_{2}(v)$ of Lazy-Alg$(\varphi,k)$ is a $(1-\varepsilon)$-covering of $B_{2}(v)$ if the following holds: i) $\hat{B}_{2}(v)\subseteq B_{2}(v)$; ii) $\mathbb{E}\left[|\hat{B}_{2}(v)|\right]\geq(1-\varepsilon)|B_{2}(v)|$, where expectation is taken over the randomness of the algorithm and/or the input sequence. When the algorithm produces a $(1-\varepsilon)$-covering $\hat{B}_{2}(v)$ of $B_{2}(v)$ for every $v$, we say it has approximation ratio $\frac{1}{(1-\varepsilon)}$.
Our main result in this section is formalized in the following
Theorem 2.
Let $\varepsilon\in(0,1)$, and fix parameters $k=0$ and $\varphi=\frac{\varepsilon}{1-\varepsilon}$. Consider any graph $G(V,E)$ submitted as a uniform random permutation of its edge set $E$ to Lazy-Alg$(\varphi,k)$. Then, at every time step $t\leq|E|$, the algorithm has approximation ratio $\frac{1}{1-\varepsilon}$. Moreover, for every $\alpha>0$ and every vertex $v\in V$, we have:
(1) $\mathbb{P}\left(|\hat{B}_{2}(v)|<\frac{1-\alpha}{1+\varphi}|L_{2}(v)|\right)\leq e^{-\frac{2\alpha^{2}|L_{2}(v)|}{(1+\varphi)^{2}}}.$
Proof.
Fix a vertex $v\in V$ and a round $t\geq 1$. In the remainder of this proof, all quantities are taken at time $t$. We are interested in how close $|\hat{B}_{2}(v)|$ is to $|B_{2}(v)|$. To begin, we note that the following relationship holds deterministically:
(2) $|\hat{B}_{2}(v)|=1+|L_{1}(v)|+|\hat{B}_{2}(v)\cap L_{2}(v)|,$
where the only random variable on the right hand side is the last term.
Figure 2. Example of a partition of $L_{2}(v)$ into three sets $C_{1},C_{2},C_{3}$. Edges connecting vertices $w\in L_{2}(v)$ to their respective partitions are thicker.
We next define a partition $\mathcal{C}=\{C_{u}:u\in L_{1}(v)\}$ of $L_{2}(v)$ as follows: for each $w\in L_{2}(v)$, we choose a vertex $u\in L_{1}(v)\cap\mathcal{N}(w)$ and assign $w$ to $C_{u}$. This way, each vertex $w\in L_{2}(v)$ is associated to exactly one edge connecting one vertex in $L_{1}(v)$ to $w$ (see Figure 2, where the edges in question are thick in the picture). Let $E_{v}$ denote the set of such edges and note that i) $E_{v}$ is a subset of the edges connecting vertices in $L_{1}(v)$ to those in $L_{2}(v)$, ii) $|E_{v}|=|L_{2}(v)|$ by definition and iii) $|C_{u}|\leq\deg_{u}-1$ for every $u\in L_{1}(v)$, given that $C_{u}$ contains a subset of $u$’s neighbors and $(v,u)$ is always present. Moreover, Algorithm 3 guarantees that $|\hat{B}_{2}(v)\cap L_{2}(v)|$ is at least the number of edges in $E_{v}$ that are black. These considerations allow us to conclude that
$|\hat{B}_{2}(v)\cap L_{2}(v)|\geq|\{e\in E_{v}:\text{ $e$ is black}\}|.$
A key observation at this point is that Algorithm 3 implies that for every $x\in V$, $\delta_{x}\leq\left\lfloor\frac{\varphi}{1+\varphi}\deg_{x}\right\rfloor$. As a consequence, if some $e=(u,w)\in E_{v}$ was not among the last $\left\lfloor\frac{\varphi}{1+\varphi}\deg_{u}\right\rfloor$ edges incident in $u$ that were released within time $t$, it is necessarily black. For $e=(u,w)\in E_{v}$, with $u\in L_{1}(v)$ and $w\in L_{2}(v)$, let $X_{e}=1$ if $e$ was among the first $\deg_{u}-\left\lfloor\frac{\varphi}{1+\varphi}\deg_{u}\right\rfloor$ edges incident in $u$ that were released up to time $t$ and let $X_{e}=0$ otherwise. Following the argument above, the event $(X_{e}=1)$ implies the event "$e$ is black", whence:
(3) $|\hat{B}_{2}(v)\cap L_{2}(v)|\geq|\{e\in E_{v}:\text{ $e$ is black}\}|\geq\sum_{e\in E_{v}}X_{e}.$
Next, we are interested in bounds on $\mathbb{P}\left(X_{e}=1\right)$. Assume $e$ is incident in $u$ and let $S$ be the set of edges incident in $u$ observed up to time $t$. Then, from 1, the sequence in which these edges are observed is just a random permutation of $S$. This immediately implies that, if $e$ is incident to vertex $u\in L_{1}(v)$, then
$\mathbb{P}\left(X_{e}=1\right)=\frac{\deg_{u}-\left\lfloor\frac{\varphi}{1+\varphi}\deg_{u}\right\rfloor}{\deg_{u}}\geq\frac{1}{1+\varphi}.$
Together with (3) this yields:
$\mathbb{E}\left[|\hat{B}_{2}(v)|\right]\geq 1+|L_{1}(v)|+\frac{1}{1+\varphi}|L_{2}(v)|\geq\frac{1}{1+\varphi}|B_{2}(v)|.$
We next show that $\sum_{e\in E_{v}}X_{e}$ is concentrated around its expectation when $|L_{2}(v)|$ is large enough, which implies that $|\hat{B}_{2}(v)|$ is concentrated around a value close to $|B_{2}(v)|$ in this case. The main technical hurdle here is that the $X_{e}$’s are correlated (albeit mildly, as we shall see). To prove concentration, we resort to Martingale properties of random edge sequences to apply the method of (Average) Bounded Differences (dubhashi2009concentration, ). In order to do this, we need bounds on $\mathbb{P}\left(X_{e}=1|X_{f}=1\right)$ and $\mathbb{P}\left(X_{e}=1|X_{f}=0\right)$, for $e,f\in E_{v}$. Assume again that $e$ is incident in $u\in L_{1}(v)$ and that $S$ is the set of edges incident in $u$ observed up to time $t$. Assume first that $f$ is also incident in $u$ and that, without loss of generality, $f$ is the $i$-th edge to appear among those in $S$. $X_{f}=1$ implies $i\leq\deg_{u}-\left\lfloor\frac{\varphi}{1+\varphi}\deg_{u}\right\rfloor$. On the other hand, for any such choice for $f$’s position in the sequence, Fact 1 implies that $e$ will appear in a position $j$ that is sampled uniformly at random from the remaining ones, so that $\mathbb{P}\left(X_{e}=1|X_{f}=1\right)=\frac{\deg_{u}-\left\lfloor\frac{\varphi}{1+\varphi}\deg_{u}\right\rfloor-1}{\deg_{u}-1}$ in this case. With a similar argument, it can be seen that $\mathbb{P}\left(X_{e}=1|X_{f}=0\right)=\frac{\deg_{u}-\left\lfloor\frac{\varphi}{1+\varphi}\deg_{u}\right\rfloor}{\deg_{u}-1}$. Intuitively and unsurprisingly, the events $(X_{e}=1)$ and $(X_{f}=1)$ are negatively correlated, while $(X_{e}=1)$ and $(X_{f}=0)$ are positively correlated. This allows us to conclude that $\mathbb{P}\left(X_{e}=1|X_{f}=1\right)\leq\mathbb{P}\left(X_{e}=1|X_{f}=0\right)$ and
$\mathbb{P}\left(X_{e}=1|X_{f}=0\right)-\mathbb{P}\left(X_{e}=1|X_{f}=1\right)\leq\frac{1}{\deg_{u}-1}.$
Assume next that $f$ is not incident in $u$. Again from Fact 1, in this case $f$ has no bearing on the relative order in which edges incident in $u$ appear, so that $\mathbb{P}\left(X_{e}=1|X_{f}=0\right)=\mathbb{P}\left(X_{e}=1|X_{f}=1\right)=\mathbb{P}\left(X_{e}=1\right)$. Now, without loss of generality, suppose that $f=(z,w)$, with $z\in L_{1}(v)$, so that $w\in C_{z}$. Denote by $E_{v}(z)$ the subset of edges in $E_{v}$ with one end point in $C_{z}$. Moving to conditional expectations we have
$\mathbb{E}\left[\sum_{e\in E_{v}}X_{e}\,|\,X_{f}=0\right]-\mathbb{E}\left[\sum_{e\in E_{v}}X_{e}\,|\,X_{f}=1\right]$
$=\sum_{e\in E_{v}}\left(\mathbb{P}\left(X_{e}=1\,|\,X_{f}=0\right)-\mathbb{P}\left(X_{e}=1\,|\,X_{f}=1\right)\right)$
$=\sum_{e\in E_{v}\setminus E_{v}(z)}\left(\mathbb{P}\left(X_{e}=1\,|\,X_{f}=0\right)-\mathbb{P}\left(X_{e}=1\,|\,X_{f}=1\right)\right)$
$+\sum_{e\in E_{v}(z)}\left(\mathbb{P}\left(X_{e}=1\,|\,X_{f}=0\right)-\mathbb{P}\left(X_{e}=1\,|\,X_{f}=1\right)\right)$
$\leq\frac{|C_{z}|}{\deg_{z}-1}\leq 1,$
where the third inequality follows from the definition of $C_{z}$, since $f$ is incident in $z$, while the last inequality follows since $|C_{z}|\leq\deg_{z}-1$ for every $z\in L_{1}(v)$, because one of the edges incident in $z$ is by definition the one shared with $v$.
We can therefore apply (dubhashi2009concentration, , Definition 5.5 and Corollary 5.1), with $c\leq|L_{2}(v)|$ to obtain, for every $\alpha>0$:
$\mathbb{P}\left(\mathbb{E}\left[\sum_{e\in E_{v}}X_{e}\right]-\sum_{e\in E_{v}}X_{e}>\alpha\mathbb{E}\left[\sum_{e\in E_{v}}X_{e}\right]\right)\leq e^{-\frac{2\alpha^{2}|L_{2}(v)|}{(1+\varphi)^{2}}},$
where in the right hand side we also used the bound $\mathbb{E}\left[\sum_{e\in E_{v}}X_{e}\right]\geq\frac{1}{1+\varphi}|L_{2}(v)|$ we showed earlier. Finally, we recall (2) and (3) to conclude that $|\hat{B}_{2}(v)|\geq\frac{1-\alpha}{1+\varphi}|L_{2}(v)|$ with (at least) the same probability. ∎
Theorem 2 easily implies approximation bounds on indices that depend on the union and/or intersection of $2$-balls. For example, we immediately have the following approximation bound on the Jaccard similarity between any pair of $2$-balls.
Corollary 3.
Under the same assumptions as Theorem 2, at any time step $t\geq 1$ and for any pair of vertices $u,v\in V$, Lazy-Alg$(\varphi,k)$ guarantees the following approximation of the Jaccard similarity between $B_{2}(u)$ and $B_{2}(v)$ with probability at least $1-e^{-\frac{2\alpha^{2}|L_{2}(u)|}{(1+\varphi)^{2}}}-e^{-\frac{2\alpha^{2}|L_{2}(v)|}{(1+\varphi)^{2}}}$:
(4) $\dfrac{\mathbf{J}(B_{2}(u),B_{2}(v))}{1-2\varepsilon^{\prime}}\geq\mathbf{J}(\hat{B}_{2}(u),\hat{B}_{2}(v))\geq(1-\varepsilon^{\prime})\mathbf{J}(B_{2}(u),B_{2}(v))-\varepsilon^{\prime},$
where $\varepsilon^{\prime}=\frac{\varphi+\alpha}{1+\varphi}$.
Proof.
It is easy to see that $|\hat{B}_{2}(u)|\geq(1-\varepsilon^{\prime})|B_{2}(u)|$ and $|\hat{B}_{2}(v)|\geq(1-\varepsilon^{\prime})|B_{2}(v)|$ together imply (4) deterministically. The result then immediately follows from Theorem 2 and a union bound on the events $(|\hat{B}_{2}(u)|<(1-\varepsilon^{\prime})|L_{2}(u)|)$ and $(|\hat{B}_{2}(v)|<(1-\varepsilon^{\prime})|L_{2}(v)|)$. ∎
4. Adversarial edge sequences
[…]
4.1. A lower bound for adversarial sequences
The lower bound for the adversarial framework described above is formalized in the following result on the approximation ratio (see Def. 1)
Theorem 1.
For every $\varphi\in[0,1]$, and integer $k\geq 0$, if Lazy-Alg$(\varphi,k)$ has approximation ratio $\rho\geq 1$, then it must have an amortized update cost of $\Omega(\Delta/\rho^{3})$, where $\Delta$ is the maximum degree of the graph.
Figure 3. Black edges are present at $t=0$, while red ones are inserted in the interval $\{1,2,\ldots,\Delta\rho^{2}\}$. At time $t>0$, an edge with one endpoint in $u_{1t\mod\Delta}$ and the other in a distinct 0-degree vertex in $S_{2}$ is added.
Proof.
Fix $\rho\geq 1$, and assume that Lazy-Alg$(\varphi,k)$ has an approximation ratio of at most $\rho$. We will show that there exist an initial graph $G^{(0)}$ with degree $\Delta$ and a sequence of edge insertions against which Lazy-Alg$(\varphi,k)$ must incur an amortized update time of $\Omega(\Delta/\rho^{3})$.
Note that for $k>0$, the algorithm is randomized. In order to address this, we prove our lower bound for every possible realization of the randomness used by the algorithm. Therefore, we assume the values of the random bits used by Lazy-Alg$(\varphi,k)$ are fixed arbitrarily (and optimally) and we assume henceforth that the behavior of the algorithm is completely deterministic.
The initial graph $G^{(0)}$ consists on $2\Delta$ vertices forming a complete bipartite graph with sides $S_{0}=\{u_{01},\dots,u_{0\Delta}\}$ and $S_{1}=\{u_{11},\dots,u_{1\Delta}\}$, along with an additional set $S_{2}$ of $\Delta\rho^{2}$ isolated vertices (see Figure 3). The sequence of edge insertions is defined as follows: for each vertex in $S_{1}$, we insert $\rho^{2}$ new edges. Each of these $\Delta\rho^{2}$ edges connects a vertex in $S_{1}$ to a previously isolated vertex in $S_{2}$.
Consider the time instant right after all edge insertions. Since we assumed that the algorithm guarantees an approximation ratio of $\rho$, it holds that for every vertex $u\in S_{0}$, $|\hat{B}_{2}(u)|\geq\frac{1}{\rho}|B_{2}(u)|=\frac{1}{\rho}(2\Delta+\Delta\rho^{2})=\Delta\rho+2\Delta/\rho$. This implies that after all edge insertions, $u$ must be aware of at least $\Delta\rho+2\Delta/\rho-2\Delta=\Delta\rho-2\Delta(1-1/\rho)$ vertices from $S_{2}$.
We say that there is a message from $v$ to $u$ if vertex $v$ performs a union operation of the form $\hat{B}_{2}(u)\leftarrow\hat{B}_{2}(u)\cup\hat{B}_{1}(v)$.
Since, at any time, every vertex $v\in S_{1}$ is adjacent to at most $\rho^{2}$ vertices in $S_{2}$, it must be that each $u\in S_{0}$ must have received at least $\frac{\Delta\rho-2\Delta(1-1/\rho)}{\rho^{2}}=\Omega(\Delta/\rho)$ messages from vertices in $S_{1}$. As a consequence, the total number of messages are at least $\Omega(\Delta^{2}/\rho)$. As the number of insertions is $\Delta\rho^{2}$, the amortized update cost per insertion is at least $\frac{\Omega(\Delta^{2}/\rho)}{\Delta\rho^{2}}=\Omega(\Delta/\rho^{3})$. ∎
We have special cases as corollaries. We need amortized update cost $\Omega(\Delta)$ if we want $\rho=O(1)$, $\Omega(\sqrt[4]{\Delta})$ if we want $\rho=O(\sqrt[4]{\Delta})$ and so on.
Remark 1.
The lower bound in Theorem 1 in fact holds for a wider class of algorithms. Informally speaking, this class includes any local algorithm that limits its online updates to the 2-hop neighbors of $u$ and $v$ only. Making this claim more formal requires addressing several technical issues that are outside the scope of the present work.
4.2. Locally $\gamma$-sparse graphs
In this section, we provide the characterization of a class of graphs for which our lazy-update approach always guarantees good amortized cost/approximation trade-offs, even under the assumption of adversarial edge insertion sequences.
Given a graph $G(V,E)$ and a subset $V^{\prime}\subseteq V$, we denote by $G[V^{\prime}]$ the subgraph induced by $V^{\prime}$. Informally, a graph is locally $\gamma$-sparse if every node in $B_{2}(u)\setminus\{u\}$ has roughly at most $\gamma$ neighbors in $L_{1}(u)$. This notion can be formalized as follows.
Definition 4.0 (locally $\gamma$-sparse graphs).
Let $\gamma\in\{0,1,\ldots,n-2\}$. A graph $G(V,E)$ is said locally $\gamma$-sparse if for each vertex $u\in V$: (i) $\forall v\in L_{1}(u)$ the degree of $v$ in $G[L_{1}(u)]$ is at most $\gamma$, and (ii) $\forall w\in L_{2}(u)$ the degree of $w$ in $G[L_{1}(v)\cup\{w\}]$ is at most $\gamma+1$.
Observe that the class of locally $\gamma$-sparse graphs grows monotonically with $\gamma$, including all possible graphs for $\gamma=n-2$, while the most restricted class is obtained for $\gamma=0$. It is interesting to note that locally $\gamma$-sparse graphs are not necessarily sparse in absolute terms. For example, for $\gamma=0$, the class coincides with the well-known class of graphs with girth at least $5$: these graphs can have up to $\Theta(n^{\frac{3}{2}})$ edges assuming Erdös’ Girth Conjecture (erdos1965some, ) (the proof of such equivalence is given in Section A.2).
A first, preliminary analysis of our lazy-update approach considers the deterministic version of Algorithm 3, i.e., when $k=0$. It turns out that this version achieves an approximation ratio of $\frac{\gamma+1}{1-\varepsilon}$ and amortized cost $O(1/\varepsilon)$ (see Section A.2). So, the approximation accuracy decreases linearly in the parameter $\gamma$. Interestingly enough, we instead show that a suitable number of random light updates allows Algorithm 3 to perform much better than its deterministic version. This is the main result of this section and it is formalized in the following theorem.
Theorem 3.
Let $\varepsilon\in(0,1)$, and let $G^{(0)}$ be an initial graph. Consider any sequence of edge insertions that yields a final graph $G$. If $G$ is locally $\gamma$-sparse  Lazy-Alg$\left(\varphi=1,nk=\frac{4(\gamma+1)}{\varepsilon}\right)$ has approximation ratio of $\frac{1}{1-\varepsilon}$ and amortized cost $O\left(\frac{\gamma+1}{\varepsilon}\right)$.
Proof of Theorem 3
For the remainder of this proof, we define the notion of quasi-black edge. A red edge $(v,w)$, with $v\in L_{1}(u)$ and $w\in L_{2}(u)$, is said to be quasi-black for $u$ if $u$ has been randomly selected by $v$ at Algorithm 3 of Algorithm 3 at least once during or after the insertion of $(v,w)$, ensuring that $w\in\hat{B}_{2}(u)$.
In this section, we use ^$\hat{\ell}_{v}$ and $\ell_{v}$ to denote the number of quasi-black and red edges, respectively, that connect $v$ to vertices in $L_{2}(u)$. Similarly, we use $b_{v}$ to represent the number of black edges of $v$ having the other endpoint in $L_{2}(u)$. Notice that ^$\hat{\ell}_{v}$ is a random variable that counts how many vertices out of $\ell_{v}$ are in $\hat{B}_{2}(u)$. We now proceed by first stating a property, whose proof can be found in Section A.3.
Lemma 0.
For each $v\in L_{1}(u)$, we have $\mathbb{E}\left[\hat{\ell}_{v}\right]\geq\ell_{v}-\frac{2(b_{v}+\gamma+1)}{k}$.
Now, let $\beta$ denote the number of vertices in $L_{2}(u)$ that have at least one black edge from $L_{1}(u)$. Consequently, these vertices are included in $\hat{B}_{2}(u)$. We have the following:
Lemma 0.
Let $G=(V,E)$ be locally $\gamma$-sparse, and $u\in V$. Then
$\beta\geq\sum_{v\in L_{1}(u)}\frac{b_{v}}{\gamma+1}.$
Proof.
The inequality follows from the fact that every node in $L_{2}(u)$ can have at most $\gamma+1$ neighbors in $L_{1}(u)$. ∎
We are now ready to prove Theorem 3.
The amortized update cost follows directly from Lemma 1.
For the approximation quality, let us consider any vertex $u\in V$. For technical convenience, we will define a subgraph $\widetilde{G}$ of $G$ by removing suitable edges from $G$, and we establish the following two properties: (i) if $k\geq\frac{2(\gamma+1)}{\varepsilon}$, the Lazy-Alg guarantees a $(1-\varepsilon)$-covering of $B_{2}(u)$ when the sequence of edge insertion is restricted to edges in $\widetilde{G}$; (ii) property (i) implies that the Lazy-Alg also guarantees a $(1-\varepsilon)$-covering of $B_{2}(u)$ for $G$, provided that $k\geq\frac{4(\gamma+1)}{\varepsilon}$.
The subgraph $\widetilde{G}$ is obtained from $G$ through the following process. For each vertex $w\in L_{2}(u)$, if there exists a black edge $(v,w)$ with $v\in L_{1}(u)$, then we remove all the red edges incident to $w$ that originate from $L_{1}(u)$. Otherwise, if all edges from $L_{1}(u)$ to $w$ are red, we retain only one and remove the rest (see Figure 4).
Figure 4. The $2$-hop neighborhood of a vertex $u$ (left), and its corresponding structure in the subgraph $\widetilde{G}$ (right).
We now prove property (i). We analyze the process at a generic time $t>0$. We want to prove that $\mathbb{E}\left[|\hat{B}_{2}(u)|\right]\geq(1-\varepsilon)|B_{2}(u)|$, for any vertex $u\in V$. Since $L_{1}(u)$ is always included in $\hat{B}_{2}(u)$, it is sufficient to prove that $|\hat{B}_{2}(u)\cap L_{2}(u)|\geq(1-\varepsilon)|L_{2}(u)|$ in expectation.
By construction of $\widetilde{G}$, we have that $|\hat{B}_{2}(u)\cap L_{2}(u)|=\beta+\sum_{v\in L_{1}(v)}\hat{\ell}_{v}$, while $|L_{2}(u)|=\beta+\sum_{v\in L_{1}(v)}\ell_{v}$. Thus, we want to show that
(5) $\sum_{v\in L_{1}(v)}\hat{\ell}_{v}\geq(1-\varepsilon)\sum_{v\in L_{1}(v)}\ell_{v}-\varepsilon\beta.$
By Lemma 5, (5) is true when
(6) $\sum_{v\in L_{1}(u)}{\hat{\ell}_{v}}$ $\geq(1-\varepsilon)\sum_{v\in L_{1}(u)}{\ell_{v}}-\frac{\varepsilon}{\gamma+1}\sum_{v\in L_{1}(u)}{b_{v}}$
$=\sum_{v\in L_{1}(u)}{\left((1-\varepsilon)\ell_{v}-\frac{\varepsilon}{\gamma+1}b_{v}\right)}.$
In turn, the inequality in (6) holds in expectation if it holds term-by-term, i.e., when
$\mathbb{E}\left[\hat{\ell}_{v}\right]\geq(1-\varepsilon)\ell_{v}-\frac{\varepsilon}{1+\gamma}b_{v},\;\;\forall v\in L_{1}(u).$
Clearly, if $\ell_{v}=0$ then ^$\hat{\ell}_{v}=0$ and the inequality holds; thus we focus on the case $\ell_{v}>0$. From Lemma 4 we know that $\mathbb{E}\left[\hat{\ell}_{v}\right]\geq\ell_{v}-\frac{2(b_{v}+\gamma+1)}{k}$. By setting $k\geq\frac{2(\gamma+1)}{\varepsilon}$ we have
$\ell_{v}-\frac{2(b_{v}+\gamma+1)}{k}\geq\ell_{v}-\frac{\varepsilon}{\gamma+1}b_{v}\geq(1-\varepsilon)\ell_{v}-\frac{\varepsilon}{\gamma+1}b_{v},$
therefore proving property (i).
Finally, we prove (ii). Notice that we build $\widetilde{G}$ by removing only red edges from $G$. Since $\varphi=1$, the degrees of the vertices $v\in L_{1}(u)$ in $\widetilde{G}$ are at most halved compared to $G$. As a consequence, the probability that $v$ selects $u$ at Algorithm 3 of Algorithm 3 in $G$ is at most half of the probability we have in $\widetilde{G}$. Therefore, doubling the value of $k$ to $\frac{4(\gamma+1)}{\varepsilon}$ ensures that the analysis we conducted on $\widetilde{G}$ also holds on $G$, guaranteeing a $(1-\varepsilon)$-covering of $B_{2}(u)$ in expectation on $G$ as well.
5. Experimental analysis
The purpose of this experimental analysis, conducted on real datasets, is the validation of our lazy approach along three main axes: i) the qualitative consistency between the theoretical findings from Sections 2, 3 and 4 and the actual behavior of our lazy approach on real, incremental datasets; ii) the accuracy of our approach on two key neighborhood-based queries, namely, size and Jaccard similarity; iii) its computational savings with respect to non-lazy baselines on medium and large incremental graphs. An extensive and complete set of results is reported in Appendix B.
5.1. Experimental setup
Platform.
Our experiments were performed on a machine with 2.3 GHz Intel Xeon Gold 5118 CPU with 24 cores, 192 GB of RAM, cache L1 32KB, shared L3 of 16MB and UMA architecture. The whole code is written in C++, compiled with GCC 10 and with the following compilation flags: -DARCH__\__X86__\__64 -Wall -Wextra -g -pg -O3 -lm. [^12] Code available at https://github.com/Gnumlab/graph_ball.
Algorithms.
[…]
i) the exact algorithm 1, which is not scalable and is only used in the first round of experiments on smaller datasets, in order to isolate the error introduced by lazy updates;
ii) the naive sketch-based baseline, which adopts sketch-based representations of $1$- and $2$-balls, but performs all necessary updates. It corresponds to Algorithm 3 with $\varphi=0$ and $k=0$, but where $1$- and/or $2$-ball unions correspond to merging the corresponding sketches. [^13] Again, the specific sketch (or sketches) used depends on the neighborhood queries one wants to support.
With the exception of the first set of experiments (i.e. item (i) above), we compared our algorithm to the sketch-based baseline. For some results, we needed to compute the true value of the parameter of interest for $2$-balls (e.g., size) by executing a suitably optimized BFS. For our Lazy-Alg, we used combinations of the following values: $\varphi=0.1,0.25,0.5,0.75,1$ and $k=0,2,4,8$.
Implementation details.
To best assess the performance of algorithms, it would be ideal to minimize the overhead deriving from the management of edge insertions in the graph. We observe that this overhead is the same for all algorithms we tested. Hence, we represented graphs using the compressed sparse row format (Eisenstat1982YaleSM, ) and, since we knew the edge insertion sequence in advance, we pre-allocated the memory needed to accommodate them, so as to minimize overhead. For experiments that required hash functions, we used tabulation hashing (tabulation-hashing, ).
Datasets.
We considered real incremental graphs of different sizes, both directed and undirected, available from NetworkRepository (networkrepository, ). In our time analysis, we also extracted a large incremental dataset from a large static graph, namely soc-friendster (soc-friendster, ), available from SNAP (snapnets, ). Following previous work on dynamic graphs (chen2022dynamictree, ; Hanauer22DynamicSurvey, ), we generated an incremental graph by adding edges sequentially and in random order, starting from an empty graph. The main features of these datasets are summarized in Table 1.
Dataset $|V|$ $|E|$ Directed Dynamic
comm-linux-kernel-reply 27,927 242,976
fb-wosn-friends (fb-wosn-friends, ) 63,731 817,090
ia-enron-email-all 87,273 321,918
soc-flickr-growth 2,302,925 33,140,017
soc-youtube-growth 3,223,589 9,376,594
soc-friendster (soc-friendster, ) 65,608,366 1,806,067,135
Table 1. Summary table of real networks used in the experiments.
5.2. Results
As we remarked in the introduction, we focused on the case of undirected graphs for ease of exposition but our approach extends seamlessly to directed graphs, such as some of the real examples we consider in this section. In such cases, the only caveat to keep in mind is that we define the $h$-ball of a vertex $u$ as the subset of vertices that are reachable from $u$ over a directed path traversing at most $h$ edges. [^14] One could alternatively define the $h$-ball of $u$ as the subset of vertices from which it is possible to reach $u$ traversing at most $h$ directed edges. We stick to the former definition, which is more frequent in social network analysis.
Impact of lazy updates.
The goal of our first experiment is twofold: i) assessing the impact of lazy updates on the estimation of $2$-balls; ii) assessing the degree of consistency between the theoretical findings of Sections 3 and 4 and the actual behavior of our algorithms on real datasets. In order to isolate the specific contribution of lazy updates in the estimation error, we implemented (true and approximate) $1$-balls and $2$-balls losslessly, as dictionaries. This way, the error in $2$-ball size estimation is only determined by our lazy update policy. Since, as we argued elsewhere in the paper, lossless representations of $2$-balls quickly becomes unfeasible for larger datasets, this first experiment was run on $3$ small-medium datasets, namely, comm-linux-kernel-reply, fb-wosn-friends and ia-enron-email-all (results for comm-linux-kernel-reply are reported in Figure 10(c) in Appendix B).
For each of the above graphs, we selected as a sample $5000$ vertices whose $2$-balls are the largest at the end of the edge insertion sequence. For every pair $(\varphi,k)$ of parameter values for Lazy-Alg$(\varphi,k)$, we performed $10$ independent runs. Each run is organized into the following steps: 1) the initial graph $G^{(0)}(V,E^{(0)})$ corresponds to the first $20\%$ edge insertions; 2) we measure the coverage (see Definition 1) of each of the $5000$ $2$-balls above by Lazy-Alg$(\varphi,k)$ at each of the $100$ equally spaced timestamps, the same in each run. For the generic timestamp $t$, we measure the average coverage
$C_{t}=\frac{1}{5000}\sum_{i=1}^{5000}\frac{\hat{B}_{i}}{B_{i}},$
where $B_{i}$ and $\hat{B}_{i}$ respectively denote the true and estimated sizes of the $i$-th $2$-ball from the sample. Finally, for each timestamp $t$, we plot the average of the $10$ values of $C_{t}$ computed in every run. The results, summarized in Figure 5, are fully consistent with our theoretical findings from Section 3. At least for the diverse dataset sample considered here, uniform random permutations are a reasonable theoretical proxy of real sequences. More in general, real sequences seem to be rather far from the pathological worst-cases analyzed in Section 4.1, so that the actual behavior of our algorithm is not only in line, but better than our analysis predicts. We also have an initial insight into the effects of the parameters $\varphi$ and $k$, which will be examined more thoroughly in the subsequent subsections.
(a) fb-wosn-friends
(b) ia-enron-email-all
Figure 5. Average coverage $C_{t}$ from a network with $20\%$ of its edges to the end of the insertion sequence. Dashed lines show the theoretical expected coverage for random sequences (see Theorem 2).
Ball size estimation via sketches.
The goal of the next round of experiments was assessing the accuracy of our algorithms in $2$-ball size estimation, in the realistic setting in which approximate $1$- and $2$-balls are represented via state-of-art sketches that are based on probabilistic counters (trevisan/646978.711822, ). In this case, we have a compound estimation error, arising from both our lazy update policy and the use of probabilistic counters. We ran experiments on the three small-medium sized datasets considered previously, plus two large ones, namely, soc-youtube-growth and soc-flickr-growth. Again, we considered the top-$5000$ largest $2$-balls as sample. The experiments were executed as in the previous case, with the following differences: i) we considered $3$ timestamps, respectively corresponding to $50\%$, $75\%$ and $100\%$ of all edge insertions of each dataset; ii) in this case, $1$- and $2$-balls are not explicitly represented as sets (not even by the baseline). As a result, ball sizes might be overestimated and coverage has no clear meaning. We therefore use Mean Absolute Percentage Error (MAPE) to measure accuracy, defined as follows:
$\text{MAPE}=\frac{1}{5000}\sum_{i=1}^{5000}\frac{|B_{i}-\hat{B}_{i}|}{B_{i}},$
where $i$ refers to the $2$-ball size of the $i$-th sampled vertex. For each dataset, this index is computed for each of $10$ independent runs at each of the $3$ timestamps we consider. Results are summarized for all datasets and combinations $(\varphi,k)$ we consider in Table 2 for the last timestamp ($100\%$ of edge insertions). Results for other timestamps ($50\%$ and $75\%$) are similar and are omitted for the sake of clarity (see Appendix B for complete results). The main takeaways here are that i) the additional error introduced by the use of sketches (which can be controlled by varying the size of the sketch) is relatively modest; ii) even relatively large values of $\varphi$ and/or small values of $k$ result in performances that are close to those of the baseline that uses sketches to represent balls, but naively performs all light updates. The effect of $k$ is more pronounced when $\varphi$ is large, contributing to a reduction in both error and, to a lesser extent, variance. Finally, we note that although the effect of increasing $k$ can be achieved by decreasing $\varphi$, our analysis in Section 4 suggests that the parameter $k$ provides robustness against worst-case scenarios. Even though the analysis of the impact of sketches on error is out of the scope of this paper, the results in the previous paragraph and the small difference with the baseline (Table 2) when using the sketches further suggest that the analysis of random sequences may also apply to real sequences.
Table 2. Mean and standard deviation of absolute percentage errors for $2$-hop neighborhood size estimation. Size estimates were made using the KMV probabilistic counter (trevisan/646978.711822, ), with size $32$. Queries were made at the end of the insertion sequence.
$k$ $\varphi=0.1$ $\varphi=0.5$ $\varphi=1$ baseline
linux 00 $0.14\pm 0.12$ $0.19\pm 0.14$ $0.17\pm 0.11$ $0.12\pm 0.10$
$2$ $0.13\pm 0.11$ $0.14\pm 0.10$ $0.16\pm 0.11$
$4$ $0.14\pm 0.09$ $0.17\pm 0.12$ $0.14\pm 0.10$
$8$ $0.14\pm 0.11$ $0.14\pm 0.09$ $0.13\pm 0.10$
fb-wosn 00 $0.16\pm 0.12$ $0.15\pm 0.11$ $0.21\pm 0.12$ $0.14\pm 0.11$
$2$ $0.13\pm 0.09$ $0.15\pm 0.10$ $0.17\pm 0.11$
$4$ $0.14\pm 0.11$ $0.15\pm 0.10$ $0.16\pm 0.11$
$8$ $0.16\pm 0.13$ $0.14\pm 0.10$ $0.15\pm 0.11$
enron 00 $0.13\pm 0.10$ $0.16\pm 0.11$ $0.20\pm 0.14$ $0.13\pm 0.11$
$2$ $0.14\pm 0.11$ $0.16\pm 0.12$ $0.16\pm 0.12$
$4$ $0.13\pm 0.12$ $0.15\pm 0.12$ $0.15\pm 0.12$
$8$ $0.13\pm 0.11$ $0.14\pm 0.10$ $0.16\pm 0.13$
flickr 00 $0.17\pm 0.14$ $0.18\pm 0.12$ $0.17\pm 0.11$ $0.17\pm 0.14$
$2$ $0.16\pm 0.12$ $0.12\pm 0.09$ $0.14\pm 0.10$
$4$ $0.14\pm 0.09$ $0.13\pm 0.10$ $0.16\pm 0.10$
$8$ $0.14\pm 0.11$ $0.13\pm 0.10$ $0.14\pm 0.09$
youtube 00 $0.15\pm 0.11$ $0.15\pm 0.10$ $0.24\pm 0.11$ $0.14\pm 0.11$
$2$ $0.16\pm 0.13$ $0.14\pm 0.10$ $0.19\pm 0.11$
$4$ $0.13\pm 0.11$ $0.13\pm 0.10$ $0.15\pm 0.11$
$8$ $0.13\pm 0.10$ $0.12\pm 0.09$ $0.15\pm 0.11$
Accuracy in Jaccard similarity estimation.
With the same goal as in the previous experiment, we now evaluate the quality of our lazy approach policy combined with the use of sketches in another graph mining task: the Jaccard similarity estimation for $2$-hop neighborhoods. The sketch used to represent the $1$- and $2$- balls is the well-known $h$-minhash (broder1997resemblance, ), with $h=100$ hash functions.
As before, $10$ independent runs were performed for each combination of parameters $\varphi,k$, and the baseline, on each of the previously used datasets. Errors were measured at $3$ different timestamps, corresponding to $50\%$, $75\%$, and $100\%$ of the edge insertion sequence.
Considering all vertex pairs in the entire graph would be computationally prohibitive, as their number is too large. Moreover, many of these pairs would have an extremely low Jaccard similarity, making it difficult to estimate them with a reasonably small error (broder2000identifying, ; broder2001completeness, ). Therefore, to better evaluate our algorithm’s and baseline’s quality, we need vertex pairs with a sufficiently high similarity in their $2$-hop neighborhoods. To address this, we adopted a similar sampling process as before: we selected the $5000$ vertices with the largest $2$-hop neighborhoods at the end of the edge insertion sequence and randomly chose $1000$ pairs whose similarity is at least $0.2$. On that sample, we evaluated the MAPE of the Jaccard similarity estimate computed using the $h$-minhash signatures. Table 3 reports the results at the end of the insertion sequence, while results for the other timestamps ($50\%$ and $75\%$) are similar and are omitted for the sake of clarity (see Appendix B for complete results).
This experiment further confirms the observations from the previous one: when using sketches to represent the $2$-balls, the errors obtained with our lazy update policy (with appropriate choices of parameters $\varphi,k$) are similar and fully comparable to those of the baseline, which performs all necessary updates.
Table 3. Mean and standard deviation of absolute percentage errors for Jaccard similarity estimation, with $100$ hash functions. Queries were made at the end of the insertion sequence.
$k$ $\varphi=0.1$ $\varphi=0.5$ $\varphi=1$ baseline
linux 00 $0.11\pm 0.09$ $0.13\pm 0.09$ $0.11\pm 0.09$ $0.10\pm 0.08$
$2$ $0.09\pm 0.07$ $0.12\pm 0.09$ $0.11\pm 0.09$
$4$ $0.10\pm 0.08$ $0.12\pm 0.09$ $0.10\pm 0.08$
$8$ $0.09\pm 0.07$ $0.12\pm 0.09$ $0.10\pm 0.08$
fb-wosn 00 $0.70\pm 0.25$ $0.72\pm 0.24$ $0.74\pm 0.23$ $0.70\pm 0.24$
$2$ $0.70\pm 0.24$ $0.70\pm 0.25$ $0.72\pm 0.24$
$4$ $0.70\pm 0.24$ $0.70\pm 0.24$ $0.72\pm 0.24$
$8$ $0.70\pm 0.24$ $0.70\pm 0.24$ $0.71\pm 0.25$
enron 00 $0.09\pm 0.09$ $0.14\pm 0.12$ $0.15\pm 0.12$ $0.09\pm 0.09$
$2$ $0.10\pm 0.09$ $0.12\pm 0.10$ $0.14\pm 0.12$
$4$ $0.09\pm 0.09$ $0.11\pm 0.10$ $0.13\pm 0.11$
$8$ $0.10\pm 0.09$ $0.10\pm 0.09$ $0.11\pm 0.10$
flickr 00 $0.12\pm 0.09$ $0.11\pm 0.09$ $0.11\pm 0.09$ $0.11\pm 0.09$
$2$ $0.11\pm 0.08$ $0.10\pm 0.08$ $0.11\pm 0.09$
$4$ $0.11\pm 0.08$ $0.10\pm 0.08$ $0.10\pm 0.08$
$8$ $0.11\pm 0.08$ $0.10\pm 0.08$ $0.11\pm 0.08$
youtube 00 $0.12\pm 0.09$ $0.11\pm 0.09$ $0.17\pm 0.13$ $0.11\pm 0.09$
$2$ $0.11\pm 0.09$ $0.11\pm 0.09$ $0.15\pm 0.11$
$4$ $0.11\pm 0.09$ $0.13\pm 0.10$ $0.16\pm 0.11$
$8$ $0.11\pm 0.09$ $0.12\pm 0.09$ $0.14\pm 0.10$
Run time analysis.
Finally, we measured the running times of our algorithm and the corresponding speed-up with respect to the baseline. In the following, we report and discuss the results for the task of size estimation, using probabilistic counters. The performances for the task of Jaccard similarity estimation are analogous, and thus reported in Table 6 in Appendix B.
Table 4 reports the average speed-up of our algorithm with respect to the naive baseline for the same combinations $(\varphi,k)$ considered previously, for all datasets except com-friendster. Speed-ups are computed in terms of total update time, i.e., the total time it takes to process the whole insertion sequence. Table 5 reports the overall processing times on com-friendster, for a subset of the combinations of $\varphi$ and $k$. It should be noted that the baseline did not complete within a reasonable amount of time in this case. Finally, Figure 6 illustrates the average time cost per operation for the baseline, as well as the slowest and fastest parameter settings of $\varphi$ and $k$ in our algorithm. These experiments clearly highlight that the number of lazy updates is the crucial factor affecting performance and that our approach is very effective at addressing this problem, resulting in considerable speed-ups that grow with the size of the graph. These results regarding processing times, in conjunction with previous findings, underscore the significant advantage in terms of speed at the expense of a modest and acceptable reduction in query quality which, due to the necessary use of sketches, is never totally accurate.
Figure 6. Average time per insertion operation (in seconds) in log-scale. The datasets are arranged on the x-axis in increasing order of number of operations, again in log-scale. The baseline time for soc-friendster dataset is not reported, since it exceeded a time limit of $36$ hrs.
Table 4. Speed up with respect to the baseline, using KMV probabilistic counters (trevisan/646978.711822, ), with size $32$.
$k$ $\varphi$
$0.1$ $0.25$ $0.5$ $0.75$ $1$
linux 0 14.39x 22.73x 30.27x 32.36x 35.65x
2 12.00x 18.22x 22.03x 21.49x 23.18x
4 11.52x 15.34x 17.63x 16.93x 17.54x
8 9.57x 11.86x 12.80x 12.03x 12.22x
fb-wosn 0 5.48x 9.40x 11.15x 15.23x 16.22x
2 5.22x 7.35x 7.16x 9.21x 9.67x
4 4.32x 6.14x 6.05x 7.18x 7.23x
8 3.79x 4.38x 4.91x 5.31x 5.22x
enron 0 4.17x 5.38x 6.64x 7.12x 7.70x
2 3.61x 4.11x 4.68x 4.82x 5.27x
4 3.05x 3.42x 3.78x 3.96x 4.21x
8 2.51x 2.73x 2.85x 3.01x 3.12x
flickr 0 13.52x 19.97x 26.54x 31.51x 34.20x
2 12.16x 16.41x 19.52x 21.93x 22.61x
4 10.76x 13.96x 16.35x 17.17x 17.79x
8 9.10x 11.19x 12.36x 12.99x 13.28x
youtube 0 41.61x 61.80x 77.08x 86.50x 93.73x
2 38.16x 51.64x 60.39x 64.93x 65.51x
4 36.08x 47.43x 52.26x 55.06x 55.90x
8 33.33x 39.60x 42.73x 43.61x 43.82x
Table 5. Total running time for ball size estimation using probabilistic counters, for com-friendster dataset. The time for the baseline algorithm is not reported since it exceeded a time limit of $36$ hrs.
$k$ $\varphi$
$0.1$ $0.25$ $0.5$ $0.75$ $1$
[…]
A quick reference for parameter setting.
The trade-off between accuracy, parameter setting and update costs can be summarized as follows: decreasing $\varphi$ increases the frequency of heavy updates (and thus accuracy), while increasing $k$ increases the number of light updates upon a single edge insertion, again improving accuracy. As our experimental results show, real edge sequences exhibit a behavior that is largely consistent with the random permutation model adopted in the analysis proposed in Section 3 and far from the adversarial sequences analyzed in Section 4. In this perspective, the choice of the parameters can be derived by our theoretical results (see Theorem 2 and Lemma 1). However, in practice, by running our algorithms on a diverse set of real networks for a multiplicity of parameter settings, we identified two robust settings, corresponding to different choices for the trade-off between accuracy and efficiency. The first setting is $\varphi=0.5$ and $k=0$, which yields a very fast solution with a reasonable accuracy. Instead, when the accuracy requirement is more stringent, we suggest setting $\varphi=0.25$ and $k=2$, which usually provides a negligible error with a reasonable speed-up.
Finally, the setting of the parameters for the sketches we used (in particular, their sizes) was simply driven by i) the trade-off established by the theoretical analyses in (trevisan/646978.711822, ) and (broder1997resemblance, ; broder2000identifying, ; CGPS24, ) between accuracy and sketch sizes and ii) the amount of memory we could afford. Ultimately, this resulted in the choices described in Section 5.1, where we used $32$-bit integers.
6. Node Centrality Ranking
In this section, we highlight the effectiveness of our approach in efficiently maintaining a ranking of the vertices of an evolving social network with respect to a popular index of centrality, a key mining task in social network analysis. In this respect, we consider the ranking corresponding to harmonic centrality (rochat2009closeness, ; boldi2014axioms, ), a crucial notion in social network analysis, quantifying the importance or influence of a vertex within a network. It is particularly valuable as a refinement of closeness centrality (freeman2002centrality, ), addressing some of its limitations (rochat2009closeness, ; boldi2014axioms, ).
Harmonic centrality.
Given a graph $G=(V,E)$, the harmonic centrality of a vertex $v\in V$ is defined as
$HC(v)=\sum_{u\in V\setminus\{v\}}\frac{1}{d_{G}(v,u)},$
where $d_{G}(v,u)$ denotes the length (i.e., number of edges) of the shortest path from $v$ to $u$ in $G$, and the value $1/d_{G}(u,v)$ is conventionally taken to be 00 when $u$ is not reachable from $v$. It is easy to see that, equivalently, harmonic centrality can be written as follows:
(7) $HC(v)=\sum_{i=1}^{n}\frac{|L_{i}(v)|}{i},$
where, $L_{i}(v)$ denotes the subset of vertices exactly $i$ hops away from $v$. Computing harmonic centrality for all vertices is expensive already in the static setting, in particular, a cost $\Omega(nm)$, corresponding to the naive approach that performs $n$ BFSs, is unavoidable in certain cases (bergamini2019computing, ). Of course, problems are magnified in the dynamic setting, where the quest for efficient algorithms, even with linear update times, is an active and promising research area (Hanauer22DynamicSurvey, ).
Maintaining harmonic centrality in evolving graphs.
From an application point of view, the primary interest is often in ranking vertices based on their harmonic centrality or maintaining the top-$k$ highest scoring vertices (bisenius2018computing, ; crescenzi2020finding, ; putman2019fast, ), rather than calculating the exact values for all of them. In this perspective our lazy approach affords accurate approximation of the actual ranking of the vertices under incremental updates and with update costs that are typically $O(1)$ on real datasets.
Our approach is straightforward: i) we approximate $HC(v)$ by truncating the summation in (7) to its first two terms; ii) rather than maintaining the sets $L_{1}(v)$ and $L_{2}(v)$, we use probabilistic counters to estimate their cardinalities, a strategy that was first explored by Boldi et al. (boldi2013core, ) in the static setting. More formally, truncated harmonic centrality is defined as
$HC_{2}(v)$ $=\sum_{u\in B_{2}(v)\setminus\{v\}}\frac{1}{d(v,u)}=|L_{1}(v)|+\frac{|L_{2}(v)|}{2}$
$=|B_{1}(v)|-1+\frac{|B_{2}(v)|-|B_{1}(v)|}{2}$
Computation of $HC_{2}(v)$ requires knowledge of the $1$-hop and $2$-hop neighborhoods of $v$ ($B_{1}(v)$ and $B_{2}(v)$ respectively). To this purpose, it is natural to approximate $HC_{2}(v)$ using the approximate sizes of $|B_{1}(v)|$ and $|B_{2}(v)|$ ($|\hat{B}_{1}(v)|$ and $|\hat{B}_{2}(v)|$, respectively), as maintained by our algorithm through probabilistic counters. Therefore, we define the approximate truncated harmonic centrality $\widetilde{HC}_{2}(v)$, an estimator of $HC_{2}(v)$, as follows
(8) $HC_{2}(v)\approx\widetilde{HC}_{2}(v)=|\hat{B}_{1}(v)|-1+\frac{|\hat{B}_{2}(v)|-|\hat{B}_{1}(v)|}{2}.$
In the remainder of this section, we show that vertex ranking obtained using $HC(v)$ and $HC_{2}(v)$ are strongly correlated, justifying our approach to a task that would otherwise be hard to solve in practice for medium or large networks. This shows that our approach enables accurate approximations of important graph properties in dynamic settings, without incurring the overhead of full neighborhood tracking.
Empirical analysis.
We perform two experiments to evaluate the effectiveness of our approximation. In both experiments, we compare three different rankings of the nodes, denoted as $\mathrm{\pi}^{*}$, $\mathrm{\pi}$ and $\widetilde{\mathrm{\pi}}$, each one computed, respectively, according to the following criteria: (i) the exact harmonic centrality, (ii) the 2-truncated harmonic centrality, and (iii) the approximated truncated harmonic centrality that makes use of our lazy approach. We use the same $3$ small-medium datasets (namely comm-linux-kernel-reply, fb-wosn-friends and, ia-enron-email-all) used in the first experiment in Section 5.2. The complexity of computing the exact harmonic centrality makes the computation on larger graphs unfeasible.
In the first experiment (Figure 7), we focus on the top-ranking nodes of $\mathrm{\pi}^{*}$, for different datasets, in different (time) snapshots (w.r.t. real edge insertion sequences). For a given value of $\alpha\in\left[0,100\right]$, we select the set $T_{\alpha}^{*}\subseteq V$ consisting of the top $\alpha\%$ nodes of $\mathrm{\pi}^{*}$. We then compute the rank correlation between $\mathrm{\pi}$ and $\mathrm{\pi}^{*}$, and between $\widetilde{\mathrm{\pi}}$ and $\mathrm{\pi}^{*}$, by restricting the rankings to the nodes in $T_{\alpha}^{*}$. This allows us to assess how well the truncated version and the corresponding approximated one preserve the relative importance of real top central nodes.
The correlations between node rankings is computed using both Spearman’s and Kendall’s $\tau$ rank correlation coefficients, while as discussed in Section 5.2, we set the algorithm parameters to $\varphi=0.25$ and $k=2$, achieving a good trade-off between update time and approximation quality [^15] Concerning the sketch, we use the KMV probabilistic counters (trevisan/646978.711822, ) with size $32$.. All reported results are averaged over $10$ independent runs, and are summarized in Figure 7. Overall, the results show that truncated harmonic centrality ($HC_{2}$) preserves a strong correlation with the exact harmonic centrality ($HC$), confirming its effectiveness as a surrogate metric. More importantly, our approximation $\widetilde{HC}_{2}$ achieves correlation scores that are only marginally lower than those of $HC_{2}$. This indicates that the approximation introduced by our method is sufficiently accurate to preserve the relative ranking, making it a good and lightweight alternative for tracking centrality in incremental networks.
In the second experiment, we evaluate how well our approach is able to recall/retrieve the top-ranked nodes of $\mathrm{\pi}^{*}$. Specifically, for different value of $\alpha\in\{1,5,10\}$, we compute the sets $T_{\alpha}^{*},T_{\alpha},\widetilde{T}_{\alpha}$ of the top $\alpha\%$ nodes of $\mathrm{\pi}^{*}$, $\mathrm{\pi}$, and $\widetilde{\mathrm{\pi}}$, respectively.
We then measure the recall of $T_{\alpha}$ and of $\widetilde{T}_{\alpha}$, defined as the fraction of nodes in $T_{\alpha}^{*}$ that also appear in $T_{\alpha}$ (resp. $\widetilde{T}_{\alpha}$), i.e., $|T_{\alpha}\cap T_{\alpha}^{*}|/|T_{\alpha}^{*}|$ (resp., $|\widetilde{T}_{\alpha}\cap T_{\alpha}^{*}|/|T_{\alpha}^{*}|$). This metric captures how effectively our algorithm identifies the truly central nodes, using only approximate $2$-hop neighborhood information.
The results are summarized in Figure 8, and report the average recall over $10$ independent runs. The reported recall values are consistently high across all datasets and at every point along the edge insertion sequence. This shows that the top nodes identified by $\widetilde{HC}_{2}$ closely match those selected by exact harmonic centrality in evolving networks. Such an accurate recovery of the most central nodes further validates the effectiveness of our algorithm.
Figure 7. Spearman’s and Kendall’s $\tau$ correlation coefficients of the $2$-truncated harmonic centrality $HC_{2}$ and its approximation $\widetilde{HC}_{2}$, with respect to the exact harmonic centrality. Each point represents the correlation of the top $\alpha\%$ vertices with respect to the ordering induced by $\pi^{*}$.
(a) Recall with $HC_{2}$
(b) Recall with $\widetilde{HC}_{2}$
Figure 8. Recall of the top $1\%$, $5\%$, and $10\%$ vertices using the $2$-truncated harmonic centrality $HC_{2}$ (on the left), and the approximate harmonic centrality $\widetilde{HC}_{2}$ (on the right). The x-axis indicate different time snapshot along the edge insertion sequence.
[…]
Final remark.
We finally note that alternative (exact or approximate) approaches not considered in this study, e.g., based on on-demand computing or approximating the quantities of interest at query time, are clearly unfeasible for (continuous) queries of the type we consider, each time requiring $n$ local computations, one for each vertex of the graph.
7. Discussion and outlook
We showed that relatively simple, lazy update algorithms can play a key role to efficiently support queries over multi-hop vertex neighborhoods on large, incremental graphs. At the same time, this work leaves a number of open questions that might deserve further investigation. A first, obvious direction is extending our approach to handle edge deletions. Actually, a simple strategy to manage a deletion of an edge $(u,v)$ is to recompute the 2-balls of the $O(\deg_{u}+\deg_{v})$ affected vertices via 2-layer BFS visits. The cost of this straightforward approach can be amortized on the entire sequence of edge updates whenever deletions are rare. However, the general case is much more challenging: while our approach is potentially useful, one of the main problems here is handling deletions when compact, sketch-based data structures are used to represent $1$- and $2$-balls. Some recent contributions address similar issues in dynamic data streams (BSS20, ; CGPS24, ), but extending our analyses to this general case does not seem straightforward. Another interesting direction is investigating strategies to handle queries over $h$-balls when $h>2$, for example maintaining their sizes under dynamic updates. In this case, each edge addition/deletion potentially has cascading effects over $h$-hops. Optimizing (amortized) update costs in this general setting does not seem trivial and we conjecture that a dependence on $h$ might be necessary. Finally, we remark that our algorithms are inherently local, i.e., whether or not to perform updates involving any vertex $v$ only depends on $v$’s immediate neighborhood. As a consequence, a potentially interesting avenue for further research is to investigate distributed variants of our approach, possibly with massive parallel architectures in mind (lee2012parallel, ).
Appendix A Proofs from Section 4
A.1. On the girth of locally $\gamma$-sparse graphs
Lemma 0.
Let $G=(V,E)$ be an undirected graph with girth $g(G)$. Then $G$ is locally 00-sparse if and only if $g(G)\geq 5$.
Proof.
We first prove that if $G$ is locally 00-sparse then $g(G)$ must be at least $5$. In order to prove that, we simply negate the statement and prove that if $G$ has girth $<5$ then $G$ can not be locally 00-sparse. Without loss of generality, assume that $g(G)=4$ (the case $g(G)=3$ is similar). Then there must exist a cycle $C=(u_{1},u_{2},u_{3},u_{4})$ of $4$ vertices. It is simple to see that $u_{2},u_{4}\in L_{1}(u_{1})$ and $u_{3}\in L_{2}(u_{1})$. Since $u_{3}$ is a neighbor of both $u_{2}$ and $u_{4}$, the degree of $u_{3}$ in the subgraph $G\left[L_{1}(u_{1})\cup\{u_{4}\}\right]$ is at least $2$, hence $G$ is not locally 00-sparse (see Figure 9(a)).
We now prove that if $g(G)\geq 5$ then $G$ must be locally 00-sparse. Again, we negate this statement and prove that if $G$ is not locally 00-sparse then the girth of $G$ must be less then $5$. Let us assume that $G$ is locally $\gamma$-sparse, for any $\gamma>0$, thus it is not locally 00-sparse. Since $G$ is not locally 00-sparse there exists a vertex $v\in V$ such that at least one of the following properties holds (see Figure 9(b)):
  1. (1)
    $\exists u\in L_{1}(v)$ such that the degree of $u$ in $G\left[L_{1}(v)\right]$ is greater then 00, or;
  2. (2)
    $\exists w\in L_{2}(v)$ such that the degree of $w$ in $G\left[L_{1}(v)\cup\{w\}\right]$ is greater then $1$.
In the first case, we have a cycle of $3$ vertices, then $g(G)=3$. In the second case, we have a cycle of $4$ vertices, then $g(G)=4$. In both cases $g(G)<5$. ∎
Figure 9.
A.2. Deterministic lazy-update on $\gamma$-sparse graphs
Theorem 2.
Let $\varepsilon\in(0,1)$, and let $G^{(0)}$ be an initial graph. Consider any sequence of edge insertions that yields a final graph $G$. If $G$ is locally $\gamma$-sparse, Lazy-Alg$(\varphi=\frac{\varepsilon}{1-\varepsilon},k=0)$ has an approximation ratio of $\frac{\gamma+1}{1-\varepsilon}$ and amortized update cost $O(1/\varepsilon)$.
Proof.
Recall that $\Delta_{u}$ denotes the black degree of $u$, and that Algorithm 3 guarantees that $\deg_{u}$ is at most $(1+\varphi)\Delta_{u}$. Then, it is simple to give an upper bound to the size of $B_{2}(u)$, that is $|B_{2}(u)|\leq 1+\sum_{v\in L_{1}(u)}(1+\varphi)\Delta_{v}$.Consider a vertex $v\in L_{1}(u)$. Since $G$ is locally $\gamma$-sparse, the number of neighbors of $v$ belonging to $L_{2}(u)$ is at lest $\deg_{v}-(\gamma+1)$ of which $\Delta_{v}-(\gamma+1)$ must belong to $\hat{B}_{2}(u)$. Moreover, a vertex in $L_{2}(u)$ has at most $\gamma+1$ neighbors in $L_{1}(u)$. Therefore:
$|\hat{B}_{2}(u)|$ $\geq\Delta_{u}+1+\frac{1}{\gamma+1}\sum_{v\in L_{1}(u)}(\Delta_{v}-(\gamma+1))$
$=\Delta_{u}+1+\frac{1}{\gamma+1}\sum_{v\in L_{1}(u)}\Delta_{v}-\underbrace{\frac{1}{\gamma+1}\sum_{v\in L_{1}(u)}(\gamma+1)}_{=\Delta_{u}}$
$=1+\frac{1}{\gamma+1}\sum_{v\in L_{1}(u)}\Delta_{v}.$
As a consequence, $|\hat{B}_{2}(u)|/|B_{2}(u)|\geq\frac{1}{(1+\varphi)(\gamma+1)}$. By setting $\varphi=\frac{\varepsilon}{1-\varepsilon}$, and by using Lemma 1, the claim follows. ∎
A.3. Proof of Lemma 4
Proof.
Let $e_{1},\dots,e_{\ell_{v}}$ be the red edges between $v$ and $L_{2}(u)$, and define the binary random variable ^$\hat{\ell}_{v}(i)$ that is equal to $1$ if $e_{i}$ is a quasi-black edge for $u$, 00 otherwise, for $i=1,\dots,\ell_{v}$. Thus we can express ^$\hat{\ell}_{v}=\sum_{i=1}^{\ell_{v}}\hat{\ell}_{v}(i)$, with expectation
(9) $\mathbb{E}\left[\hat{\ell}_{v}\right]$ $=\sum_{i=1}^{\ell_{v}}{\mathbb{P}\left(\hat{\ell}_{v}(i)=1\right)}=\ell_{v}-\sum_{i=1}^{\ell_{v}}{\mathbb{P}\left(\hat{\ell}_{v}(i)=0\right)}.$
Without loss of generality, assume that the edges $e_{1},\dots,e_{\ell_{v}}$ have been inserted at times $t_{1}<\dots<t_{\ell_{v}}$, respectively. If $e_{i}$ is not a quasi-black edge for $u$, then it must be that $u$ is not selected by $v$ at Algorithm 3 of Algorithm 3, at times $t_{i},t_{i+1},\dots,t_{\ell_{v}}$. This holds with probability
(10) (^$\mathbb{P}\left(\hat{\ell}_{v}(i)=0\right)\leq\prod_{j=i}^{\ell_{v}}\left(1-\frac{k}{\deg_{v}^{(t_{j})}}\right)\leq\prod_{j=i}^{\ell_{v}}\left(1-\frac{k}{\deg_{v}^{(t_{\ell_{v}})}}\right)$
$\leq\left(1-\frac{k}{b_{v}+\ell_{v}+\gamma+1}\right)^{\ell_{v}-i+1}\leq\left(1-\frac{k}{2(b_{v}+\gamma+1)}\right)^{\ell_{v}-i}.$
The third inequality holds since the edges incident to $v$ having endpoints in $L_{1}(u)$ are at most $\gamma$, while those having endpoints in $L_{2}(u)$ are exactly $b_{v}+\ell_{v}$. Moreover, the last inequality holds because $\ell_{v}\leq\delta_{v}\leq\Delta_{v}\leq b_{v}+\gamma+1$, given the assumption $\varphi=1$.
By plugging in (10) into (9) and we obtain
$\mathbb{E}\left[\hat{\ell}_{v}\right]\geq\ell_{v}-\sum_{i=1}^{\ell_{v}}\left(1-\frac{k}{2(b_{v}+\gamma+1)}\right)^{\ell_{v}-i}$
$=\ell_{v}-\sum_{i=0}^{\ell_{v}-1}\left(1-\frac{k}{2(b_{v}+\gamma+1)}\right)^{i}\leq\ell_{v}-\frac{1-\left(1-\frac{k}{2(b_{v}+\gamma+1)}\right)^{\ell_{v}}}{1-\left(1-\frac{k}{2(b_{v}+\gamma+1)}\right)}$
$\geq\ell_{v}-\frac{1}{1-\left(1-\frac{k}{2(b_{v}+\gamma+1)}\right)}\geq\ell_{v}-\frac{2(b_{v}+\gamma+1)}{k}.$
Appendix B Additional material for Section 5
In this section, we present the comprehensive set of experimental results, providing an exhaustive overview and offering a detailed picture that completes the material available in Section 5.
(a) fb-wosn-friends
(b) ia-enron-email-all
(c) comm-linux-kernel-reply
Figure 10. Average coverage $C_{t}$ from a network with $20\%$ of its edges to the end of the insertion sequence. Dashed lines show the theoretical expected coverage for random sequences (see Theorem 2).
Table 6. Speed up with respect to the baseline, using minhash sketches, with $100$ hash functions.
$k$ $\varphi$
$0.1$ $0.25$ $0.5$ $0.75$ $1$
linux 0 6.73x 8.01x 8.67x 8.96x 9.08x
2 6.35x 7.36x 7.85x 8.03x 8.11x
4 6.04x 6.90x 7.26x 7.38x 7.45x
8 5.52x 6.13x 6.35x 6.40x 6.45x
fb-wosn 0 2.55x 3.13x 3.47x 3.63x 3.70x
2 2.38x 2.80x 3.04x 3.13x 3.18x
4 2.27x 2.60x 2.76x 2.81x 2.86x
8 2.07x 2.28x 2.37x 2.38x 2.41x
enron 0 1.85x 1.99x 2.07x 2.10x 2.12x
2 1.77x 1.88x 1.94x 1.96x 1.97x
4 1.70x 1.80x 1.84x 1.86x 1.86x
8 1.59x 1.65x 1.67x 1.68x 1.68x
flickr 0 12.36x 17.15x 20.52x 22.13x 23.10x
2 11.18x 14.69x 16.81x 17.66x 18.19x
4 10.36x 13.11x 14.63x 15.22x 15.59x
8 9.08x 10.90x 11.76x 12.10x 12.32x
youtube 0 41.70x 53.71x 62.06x 66.81x 69.02x
2 38.74x 47.56x 52.71x 55.24x 56.28x
4 36.69x 43.71x 47.37x 49.06x 49.79x
8 33.38x 38.10x 40.17x 40.92x 41.30x
Table 7. Mean and standard deviation of absolute percentage errors for $2$-hop neighborhood size estimation. Size estimates were made using the KMV counter [5], with size $32$. Each result is done on $10$ independent runs, over a sample of the first $5000$ vertices with the largest $2$-hop neighborhood, after the insertion of $50\%$ of the edges.
$k$ $\varphi=0.1$ $\varphi=0.25$ $\varphi=0.5$ $\varphi=0.75$ $\varphi=1$ baseline
linux 00 $0.14\pm 0.12$ $0.13\pm 0.09$ $0.16\pm 0.11$ $0.21\pm 0.13$ $0.23\pm 0.13$ $0.12\pm 0.10$
$2$ $0.13\pm 0.09$ $0.15\pm 0.11$ $0.14\pm 0.10$ $0.17\pm 0.11$ $0.20\pm 0.12$
$4$ $0.13\pm 0.09$ $0.15\pm 0.11$ $0.15\pm 0.10$ $0.15\pm 0.11$ $0.17\pm 0.11$
$8$ $0.14\pm 0.11$ $0.13\pm 0.10$ $0.12\pm 0.09$ $0.20\pm 0.13$ $0.17\pm 0.11$
fb-wosn 00 $0.15\pm 0.12$ $0.16\pm 0.10$ $0.19\pm 0.11$ $0.21\pm 0.11$ $0.27\pm 0.13$ $0.13\pm 0.11$
$2$ $0.14\pm 0.10$ $0.14\pm 0.10$ $0.16\pm 0.11$ $0.20\pm 0.12$ $0.20\pm 0.11$
$4$ $0.14\pm 0.11$ $0.14\pm 0.10$ $0.15\pm 0.10$ $0.17\pm 0.12$ $0.17\pm 0.11$
$8$ $0.14\pm 0.12$ $0.14\pm 0.11$ $0.15\pm 0.11$ $0.15\pm 0.11$ $0.15\pm 0.10$
enron 00 $0.13\pm 0.10$ $0.16\pm 0.11$ $0.15\pm 0.10$ $0.18\pm 0.12$ $0.19\pm 0.13$ $0.14\pm 0.12$
$2$ $0.14\pm 0.11$ $0.14\pm 0.11$ $0.16\pm 0.12$ $0.15\pm 0.11$ $0.16\pm 0.11$
$4$ $0.14\pm 0.12$ $0.14\pm 0.11$ $0.15\pm 0.10$ $0.15\pm 0.12$ $0.15\pm 0.12$
$8$ $0.13\pm 0.10$ $0.13\pm 0.11$ $0.14\pm 0.10$ $0.15\pm 0.13$ $0.16\pm 0.13$
flickr 00 $0.19\pm 0.15$ $0.16\pm 0.12$ $0.18\pm 0.12$ $0.16\pm 0.12$ $0.17\pm 0.12$ $0.14\pm 0.12$
$2$ $0.15\pm 0.11$ $0.15\pm 0.11$ $0.12\pm 0.10$ $0.16\pm 0.13$ $0.12\pm 0.09$
$4$ $0.13\pm 0.09$ $0.15\pm 0.12$ $0.14\pm 0.11$ $0.15\pm 0.11$ $0.15\pm 0.11$
$8$ $0.13\pm 0.11$ $0.19\pm 0.15$ $0.14\pm 0.12$ $0.13\pm 0.11$ $0.13\pm 0.09$
youtube 00 $0.14\pm 0.11$ $0.16\pm 0.12$ $0.17\pm 0.11$ $0.22\pm 0.12$ $0.24\pm 0.13$ $0.15\pm 0.12$
$2$ $0.14\pm 0.10$ $0.15\pm 0.11$ $0.15\pm 0.11$ $0.18\pm 0.12$ $0.18\pm 0.11$
$4$ $0.15\pm 0.12$ $0.15\pm 0.11$ $0.15\pm 0.11$ $0.16\pm 0.11$ $0.17\pm 0.12$
$8$ $0.15\pm 0.13$ $0.14\pm 0.12$ $0.14\pm 0.11$ $0.16\pm 0.11$ $0.16\pm 0.11$
Table 8. Mean and standard deviation of absolute percentage errors for $2$-hop neighborhood size estimation. Size estimates were made using the KMV counter [5], with size $32$. Each result is done on $10$ independent runs, over a sample of the first $5000$ vertices with the largest $2$-hop neighborhood, after the insertion of $75\%$ of the edges.
$k$ $\varphi=0.1$ $\varphi=0.25$ $\varphi=0.5$ $\varphi=0.75$ $\varphi=1$ baseline
linux 00 $0.13\pm 0.11$ $0.14\pm 0.10$ $0.19\pm 0.15$ $0.18\pm 0.12$ $0.18\pm 0.12$ $0.11\pm 0.09$
$2$ $0.14\pm 0.11$ $0.15\pm 0.10$ $0.14\pm 0.10$ $0.16\pm 0.11$ $0.17\pm 0.11$
$4$ $0.14\pm 0.09$ $0.16\pm 0.11$ $0.16\pm 0.12$ $0.14\pm 0.11$ $0.15\pm 0.10$
$8$ $0.15\pm 0.13$ $0.14\pm 0.10$ $0.13\pm 0.09$ $0.19\pm 0.15$ $0.14\pm 0.10$
fb-wosn 00 $0.16\pm 0.12$ $0.15\pm 0.10$ $0.16\pm 0.11$ $0.17\pm 0.11$ $0.22\pm 0.12$ $0.14\pm 0.11$
$2$ $0.13\pm 0.10$ $0.14\pm 0.10$ $0.15\pm 0.10$ $0.17\pm 0.11$ $0.18\pm 0.11$
$4$ $0.14\pm 0.11$ $0.14\pm 0.11$ $0.15\pm 0.10$ $0.16\pm 0.13$ $0.16\pm 0.11$
$8$ $0.15\pm 0.13$ $0.13\pm 0.11$ $0.14\pm 0.11$ $0.14\pm 0.10$ $0.15\pm 0.11$
enron 00 $0.13\pm 0.10$ $0.16\pm 0.12$ $0.15\pm 0.11$ $0.18\pm 0.12$ $0.20\pm 0.13$ $0.13\pm 0.11$
$2$ $0.14\pm 0.11$ $0.14\pm 0.10$ $0.17\pm 0.12$ $0.15\pm 0.11$ $0.16\pm 0.12$
$4$ $0.14\pm 0.12$ $0.14\pm 0.10$ $0.15\pm 0.11$ $0.16\pm 0.12$ $0.16\pm 0.12$
$8$ $0.13\pm 0.10$ $0.15\pm 0.11$ $0.14\pm 0.10$ $0.14\pm 0.11$ $0.16\pm 0.13$
flickr 00 $0.18\pm 0.15$ $0.16\pm 0.11$ $0.19\pm 0.12$ $0.16\pm 0.11$ $0.18\pm 0.11$ $0.17\pm 0.14$
$2$ $0.14\pm 0.11$ $0.16\pm 0.12$ $0.12\pm 0.09$ $0.16\pm 0.11$ $0.14\pm 0.09$
$4$ $0.14\pm 0.09$ $0.17\pm 0.13$ $0.13\pm 0.10$ $0.15\pm 0.11$ $0.16\pm 0.10$
$8$ $0.13\pm 0.10$ $0.17\pm 0.13$ $0.13\pm 0.10$ $0.13\pm 0.10$ $0.14\pm 0.10$
youtube 00 $0.14\pm 0.11$ $0.15\pm 0.10$ $0.19\pm 0.11$ $0.21\pm 0.11$ $0.24\pm 0.12$ $0.13\pm 0.10$
$2$ $0.15\pm 0.13$ $0.14\pm 0.09$ $0.15\pm 0.11$ $0.15\pm 0.11$ $0.17\pm 0.10$
$4$ $0.14\pm 0.11$ $0.15\pm 0.10$ $0.16\pm 0.12$ $0.14\pm 0.09$ $0.16\pm 0.11$
$8$ $0.15\pm 0.12$ $0.14\pm 0.11$ $0.13\pm 0.09$ $0.14\pm 0.10$ $0.16\pm 0.10$
Table 9. Mean and standard deviation of absolute percentage errors for $2$-hop neighborhood size estimation. Size estimates were made using the KMV probabilistic counter [5], with size $32$. Each result is done on $10$ independent runs, over a sample of the first $5000$ vertices with the largest $2$-hop neighborhood, after the insertion of $100\%$ of the edges.
$k$ $\varphi=0.1$ $\varphi=0.25$ $\varphi=0.5$ $\varphi=0.75$ $\varphi=1$ baseline
linux 00 $0.14\pm 0.12$ $0.14\pm 0.10$ $0.19\pm 0.14$ $0.16\pm 0.11$ $0.17\pm 0.11$ $0.12\pm 0.10$
$2$ $0.13\pm 0.11$ $0.16\pm 0.10$ $0.14\pm 0.10$ $0.15\pm 0.11$ $0.16\pm 0.11$
$4$ $0.14\pm 0.09$ $0.16\pm 0.11$ $0.17\pm 0.12$ $0.15\pm 0.14$ $0.14\pm 0.10$
$8$ $0.14\pm 0.11$ $0.13\pm 0.10$ $0.14\pm 0.09$ $0.21\pm 0.16$ $0.13\pm 0.10$
fb-wosn 00 $0.16\pm 0.12$ $0.15\pm 0.10$ $0.15\pm 0.11$ $0.17\pm 0.11$ $0.21\pm 0.12$ $0.14\pm 0.11$
$2$ $0.13\pm 0.09$ $0.14\pm 0.10$ $0.15\pm 0.10$ $0.17\pm 0.11$ $0.17\pm 0.11$
$4$ $0.14\pm 0.11$ $0.14\pm 0.11$ $0.15\pm 0.10$ $0.16\pm 0.12$ $0.16\pm 0.11$
$8$ $0.16\pm 0.13$ $0.13\pm 0.11$ $0.14\pm 0.10$ $0.14\pm 0.10$ $0.15\pm 0.11$
enron 00 $0.13\pm 0.10$ $0.15\pm 0.12$ $0.16\pm 0.11$ $0.18\pm 0.12$ $0.20\pm 0.14$ $0.13\pm 0.11$
$2$ $0.14\pm 0.11$ $0.14\pm 0.10$ $0.16\pm 0.12$ $0.15\pm 0.10$ $0.16\pm 0.12$
$4$ $0.13\pm 0.12$ $0.14\pm 0.10$ $0.15\pm 0.12$ $0.16\pm 0.12$ $0.15\pm 0.12$
$8$ $0.13\pm 0.11$ $0.15\pm 0.11$ $0.14\pm 0.10$ $0.14\pm 0.11$ $0.16\pm 0.13$
flickr 00 $0.17\pm 0.14$ $0.15\pm 0.11$ $0.18\pm 0.12$ $0.17\pm 0.11$ $0.17\pm 0.11$ $0.17\pm 0.14$
$2$ $0.16\pm 0.12$ $0.17\pm 0.12$ $0.12\pm 0.09$ $0.16\pm 0.11$ $0.14\pm 0.10$
$4$ $0.14\pm 0.09$ $0.16\pm 0.13$ $0.13\pm 0.10$ $0.14\pm 0.11$ $0.16\pm 0.10$
$8$ $0.14\pm 0.11$ $0.16\pm 0.12$ $0.13\pm 0.10$ $0.13\pm 0.10$ $0.14\pm 0.09$
youtube 00 $0.15\pm 0.11$ $0.13\pm 0.10$ $0.15\pm 0.10$ $0.23\pm 0.13$ $0.24\pm 0.11$ $0.14\pm 0.11$
$2$ $0.16\pm 0.13$ $0.14\pm 0.10$ $0.14\pm 0.10$ $0.15\pm 0.11$ $0.19\pm 0.11$
$4$ $0.13\pm 0.11$ $0.16\pm 0.10$ $0.13\pm 0.10$ $0.15\pm 0.09$ $0.15\pm 0.11$
$8$ $0.13\pm 0.10$ $0.15\pm 0.11$ $0.12\pm 0.09$ $0.14\pm 0.10$ $0.15\pm 0.11$
Table 10. Mean and standard deviation of absolute percentage errors for Jaccard similarity estimation, with $100$ hash functions. Queries were made after inserting the first $50\%$ of the edges.
$k$ $\varphi=0.1$ $\varphi=0.25$ $\varphi=0.5$ $\varphi=0.75$ $\varphi=1$ baseline
linux 00 $0.11\pm 0.09$ $0.11\pm 0.09$ $0.09\pm 0.08$ $0.12\pm 0.09$ $0.14\pm 0.10$ $0.08\pm 0.07$
$2$ $0.09\pm 0.07$ $0.11\pm 0.09$ $0.10\pm 0.08$ $0.11\pm 0.08$ $0.15\pm 0.11$
$4$ $0.09\pm 0.08$ $0.11\pm 0.08$ $0.11\pm 0.09$ $0.11\pm 0.09$ $0.11\pm 0.09$
$8$ $0.09\pm 0.07$ $0.11\pm 0.08$ $0.10\pm 0.08$ $0.11\pm 0.08$ $0.11\pm 0.09$
fb-wosn 00 $0.80\pm 0.25$ $0.80\pm 0.25$ $0.80\pm 0.24$ $0.81\pm 0.24$ $0.81\pm 0.24$ $0.80\pm 0.24$
$2$ $0.80\pm 0.24$ $0.79\pm 0.25$ $0.79\pm 0.25$ $0.80\pm 0.24$ $0.81\pm 0.24$
$4$ $0.79\pm 0.25$ $0.79\pm 0.25$ $0.80\pm 0.24$ $0.80\pm 0.24$ $0.80\pm 0.24$
$8$ $0.80\pm 0.24$ $0.79\pm 0.25$ $0.79\pm 0.25$ $0.80\pm 0.24$ $0.80\pm 0.24$
enron 00 $0.09\pm 0.09$ $0.11\pm 0.10$ $0.12\pm 0.10$ $0.13\pm 0.11$ $0.14\pm 0.11$ $0.09\pm 0.09$
$2$ $0.09\pm 0.09$ $0.11\pm 0.10$ $0.11\pm 0.10$ $0.12\pm 0.10$ $0.13\pm 0.11$
$4$ $0.09\pm 0.08$ $0.10\pm 0.09$ $0.11\pm 0.10$ $0.12\pm 0.10$ $0.12\pm 0.10$
$8$ $0.09\pm 0.09$ $0.10\pm 0.09$ $0.10\pm 0.09$ $0.11\pm 0.09$ $0.10\pm 0.09$
flickr 00 $0.10\pm 0.08$ $0.10\pm 0.08$ $0.11\pm 0.09$ $0.11\pm 0.09$ $0.12\pm 0.10$ $0.11\pm 0.09$
$2$ $0.10\pm 0.08$ $0.11\pm 0.09$ $0.10\pm 0.08$ $0.12\pm 0.09$ $0.11\pm 0.09$
$4$ $0.11\pm 0.08$ $0.11\pm 0.09$ $0.10\pm 0.08$ $0.11\pm 0.08$ $0.11\pm 0.08$
$8$ $0.11\pm 0.08$ $0.11\pm 0.09$ $0.10\pm 0.08$ $0.12\pm 0.09$ $0.11\pm 0.09$
youtube 00 $0.10\pm 0.09$ $0.10\pm 0.09$ $0.10\pm 0.08$ $0.11\pm 0.09$ $0.13\pm 0.11$ $0.12\pm 0.10$
$2$ $0.11\pm 0.09$ $0.11\pm 0.10$ $0.11\pm 0.09$ $0.16\pm 0.12$ $0.12\pm 0.09$
$4$ $0.10\pm 0.08$ $0.10\pm 0.08$ $0.13\pm 0.11$ $0.18\pm 0.13$ $0.13\pm 0.09$
$8$ $0.10\pm 0.08$ $0.10\pm 0.08$ $0.11\pm 0.09$ $0.15\pm 0.11$ $0.16\pm 0.11$
Table 11. Mean and standard deviation of absolute percentage errors for Jaccard similarity estimation, with $100$ hash functions. Queries were made after inserting the first $75\%$ of the edges.
$k$ $\varphi=0.1$ $\varphi=0.25$ $\varphi=0.5$ $\varphi=0.75$ $\varphi=1$ baseline
linux 00 $0.11\pm 0.09$ $0.11\pm 0.09$ $0.09\pm 0.08$ $0.12\pm 0.09$ $0.14\pm 0.10$ $0.08\pm 0.07$
$2$ $0.09\pm 0.07$ $0.11\pm 0.09$ $0.10\pm 0.08$ $0.11\pm 0.08$ $0.15\pm 0.11$
$4$ $0.09\pm 0.08$ $0.11\pm 0.08$ $0.11\pm 0.09$ $0.11\pm 0.09$ $0.11\pm 0.09$
$8$ $0.09\pm 0.07$ $0.11\pm 0.08$ $0.10\pm 0.08$ $0.11\pm 0.08$ $0.11\pm 0.09$
fb-wosn 00 $0.80\pm 0.25$ $0.80\pm 0.25$ $0.80\pm 0.24$ $0.81\pm 0.24$ $0.81\pm 0.24$ $0.80\pm 0.24$
$2$ $0.80\pm 0.24$ $0.79\pm 0.25$ $0.79\pm 0.25$ $0.80\pm 0.24$ $0.81\pm 0.24$
$4$ $0.79\pm 0.25$ $0.79\pm 0.25$ $0.80\pm 0.24$ $0.80\pm 0.24$ $0.80\pm 0.24$
$8$ $0.80\pm 0.24$ $0.79\pm 0.25$ $0.79\pm 0.25$ $0.80\pm 0.24$ $0.80\pm 0.24$
enron 00 $0.09\pm 0.09$ $0.11\pm 0.10$ $0.12\pm 0.10$ $0.13\pm 0.11$ $0.14\pm 0.11$ $0.09\pm 0.09$
$2$ $0.09\pm 0.09$ $0.11\pm 0.10$ $0.11\pm 0.10$ $0.12\pm 0.10$ $0.13\pm 0.11$
$4$ $0.09\pm 0.08$ $0.10\pm 0.09$ $0.11\pm 0.10$ $0.12\pm 0.10$ $0.12\pm 0.10$
$8$ $0.09\pm 0.09$ $0.10\pm 0.09$ $0.10\pm 0.09$ $0.11\pm 0.09$ $0.10\pm 0.09$
flickr 00 $0.10\pm 0.08$ $0.10\pm 0.08$ $0.11\pm 0.09$ $0.11\pm 0.09$ $0.12\pm 0.10$ $0.11\pm 0.09$
$2$ $0.10\pm 0.08$ $0.11\pm 0.09$ $0.10\pm 0.08$ $0.12\pm 0.09$ $0.11\pm 0.09$
$4$ $0.11\pm 0.08$ $0.11\pm 0.09$ $0.10\pm 0.08$ $0.11\pm 0.08$ $0.11\pm 0.08$
$8$ $0.11\pm 0.08$ $0.11\pm 0.09$ $0.10\pm 0.08$ $0.12\pm 0.09$ $0.11\pm 0.09$
youtube 00 $0.10\pm 0.09$ $0.10\pm 0.09$ $0.10\pm 0.08$ $0.11\pm 0.09$ $0.13\pm 0.11$ $0.12\pm 0.10$
$2$ $0.11\pm 0.09$ $0.11\pm 0.10$ $0.11\pm 0.09$ $0.16\pm 0.12$ $0.12\pm 0.09$
$4$ $0.10\pm 0.08$ $0.10\pm 0.08$ $0.13\pm 0.11$ $0.18\pm 0.13$ $0.13\pm 0.09$
$8$ $0.10\pm 0.08$ $0.10\pm 0.08$ $0.11\pm 0.09$ $0.15\pm 0.11$ $0.16\pm 0.11$
Table 12. Mean and standard deviation of absolute percentage errors for Jaccard similarity estimation, with $100$ hash functions. Queries were made after inserting the first $100\%$ of the edges.
$k$ $\varphi=0.1$ $\varphi=0.25$ $\varphi=0.5$ $\varphi=0.75$ $\varphi=1$ baseline
linux 00 $0.11\pm 0.09$ $0.11\pm 0.09$ $0.09\pm 0.08$ $0.12\pm 0.09$ $0.14\pm 0.10$ $0.08\pm 0.07$
$2$ $0.09\pm 0.07$ $0.11\pm 0.09$ $0.10\pm 0.08$ $0.11\pm 0.08$ $0.15\pm 0.11$
$4$ $0.09\pm 0.08$ $0.11\pm 0.08$ $0.11\pm 0.09$ $0.11\pm 0.09$ $0.11\pm 0.09$
$8$ $0.09\pm 0.07$ $0.11\pm 0.08$ $0.10\pm 0.08$ $0.11\pm 0.08$ $0.11\pm 0.09$
fb-wosn 00 $0.80\pm 0.25$ $0.80\pm 0.25$ $0.80\pm 0.24$ $0.81\pm 0.24$ $0.81\pm 0.24$ $0.80\pm 0.24$
$2$ $0.80\pm 0.24$ $0.79\pm 0.25$ $0.79\pm 0.25$ $0.80\pm 0.24$ $0.81\pm 0.24$
$4$ $0.79\pm 0.25$ $0.79\pm 0.25$ $0.80\pm 0.24$ $0.80\pm 0.24$ $0.80\pm 0.24$
$8$ $0.80\pm 0.24$ $0.79\pm 0.25$ $0.79\pm 0.25$ $0.80\pm 0.24$ $0.80\pm 0.24$
enron 00 $0.09\pm 0.09$ $0.11\pm 0.10$ $0.12\pm 0.10$ $0.13\pm 0.11$ $0.14\pm 0.11$ $0.09\pm 0.09$
$2$ $0.09\pm 0.09$ $0.11\pm 0.10$ $0.11\pm 0.10$ $0.12\pm 0.10$ $0.13\pm 0.11$
$4$ $0.09\pm 0.08$ $0.10\pm 0.09$ $0.11\pm 0.10$ $0.12\pm 0.10$ $0.12\pm 0.10$
$8$ $0.09\pm 0.09$ $0.10\pm 0.09$ $0.10\pm 0.09$ $0.11\pm 0.09$ $0.10\pm 0.09$
flickr 00 $0.10\pm 0.08$ $0.10\pm 0.08$ $0.11\pm 0.09$ $0.11\pm 0.09$ $0.12\pm 0.10$ $0.11\pm 0.09$
$2$ $0.10\pm 0.08$ $0.11\pm 0.09$ $0.10\pm 0.08$ $0.12\pm 0.09$ $0.11\pm 0.09$
$4$ $0.11\pm 0.08$ $0.11\pm 0.09$ $0.10\pm 0.08$ $0.11\pm 0.08$ $0.11\pm 0.08$
$8$ $0.11\pm 0.08$ $0.11\pm 0.09$ $0.10\pm 0.08$ $0.12\pm 0.09$ $0.11\pm 0.09$
youtube 00 $0.10\pm 0.09$ $0.10\pm 0.09$ $0.10\pm 0.08$ $0.11\pm 0.09$ $0.13\pm 0.11$ $0.12\pm 0.10$
$2$ $0.11\pm 0.09$ $0.11\pm 0.10$ $0.11\pm 0.09$ $0.16\pm 0.12$ $0.12\pm 0.09$
$4$ $0.10\pm 0.08$ $0.10\pm 0.08$ $0.13\pm 0.11$ $0.18\pm 0.13$ $0.13\pm 0.09$
$8$ $0.10\pm 0.08$ $0.10\pm 0.08$ $0.11\pm 0.09$ $0.15\pm 0.11$ $0.16\pm 0.11$